### v1

# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

Abstract
Recent advancements in machine learning have improved performance while also increasing computational demands. While federated and distributed setups address these issues, their structure is vulnerable to malicious influences. In this paper, we address a specific threat, Byzantine attacks, where compromised clients inject adversarial updates to derail global convergence. We combine the trust scores concept with trial function methodology to dynamically filter outliers. Our methods address the critical limitations of previous approaches, allowing functionality even when Byzantine nodes are in the majority. Moreover, our algorithms adapt to widely used scaled methods like Adam and RMSProp, as well as practical scenarios, including local training and partial participation. We validate the robustness of our methods by conducting extensive experiments on both synthetic and real ECG data collected from medical institutions. Furthermore, we provide a broad theoretical analysis of our algorithms and their extensions to aforementioned practical setups. The convergence guarantees of our methods are comparable to those of classical algorithms developed without Byzantine interference.
1  Introduction
As the field of machine learning expands, researchers are confronted with challenges stemming from more complex models, larger computational demands, and data privacy. To address these issues, distributed and federated learning scenarios are among the most popular that have been developed (Kairouz et al., 2021; Konečný et al., 2016; Li et al., 2020). These approaches are crucial for a wide range of tasks (Smith et al., 2017; McMahan et al., 2017; Verbraeken et al., 2020), yet they introduce several complications. Making learning process multi-node though leads to various threats, especially those related to data storage and transmission. These vulnerabilities can manifest as device malfunctions, incorrect data relays, or even initial data corruption (Biggio et al., 2012). Moreover, (Wang et al., 2023) demonstrated how sophisticated adversarial attacks on data integrity pose significant challenges, fundamentally compromising the effectiveness of the learning process. In the midst of them are Byzantine attacks. They occur in networks where certain workers, known as Byzantines, may deliberately or accidentally introduce corrupted data, disrupting the learning process (Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2021b). This underscores the critical need for robust defenses.
This paper specifically examines the threat of Byzantine attacks and highlights the limitations of existing protection mechanisms, particularly their reliance on strong assumptions. In response, we propose an approach that leverages diverse concepts to develop a universal method, free from these constraints and applicable to practical scenarios.
Related work. Methods resistant to Byzantine attacks are currently crucial for solving optimization problems. Classical existing methods for distributed optimization, such as SGD (Robbins & Monro, 1951; Bottou, 2012; Recht & Ré, 2013; McMahan et al., 2017), Adam (Kingma & Ba, 2014; Reddi et al., 2020) and Scaffold (Karimireddy et al., 2020), share a basic aggregation rule: averaging the gradients or models received from devices. However, they cease to operate as soon as even a single Byzantine worker appears in the network (Blanchard et al., 2017).
Given the critical importance of this problem, numerous publications addressed it (Feng et al., 2014; Damaskinos et al., 2019). Initial approaches proposed robust aggregation rules for data from devices, such as Coordinate-wise median, Trimmed Mean (Yin et al., 2018), Krum (Blanchard et al., 2017), Bulyan (Mhamdi et al., 2018). However, sophisticated attacks, such as ALIE (Baruch et al., 2019) or Inner Product Manipulation (Xie et al., 2020), managed to circumvent these aggregation rules by shifting the mean they seek to find.
Moreover, these rules are non-robust even in the absence of attacks, for example, in the case of imbalanced classes. This issue was addressed in (Karimireddy et al., 2021b), where Centered Clip (CC) technique was revealed. Additionally, the authors highlighted that the aforementioned Byzantine-robust methods cannot converge with any predetermined accuracy. Given the importance of this issue, they added client momentum, effectively tackling this problem. Another approach to combat the Byzantines is the application of the variance reduction technique, originally developed to eliminate irreducible errors in stochastic methods. It was proposed as an effective mean of mitigating the presence of noise in computing stochastic gradient estimates (Gorbunov et al., 2023). Later this idea was developed in (Malinovsky et al., 2023). These methods demonstrated significant improvements in resilience against Byzantine attacks. However, they also contain some notable limitations. First, variance reduction methods exhibit moderate convergence in deep learning applications and are prone to overfitting (Defazio & Bottou, 2019). Additionally, all aforementioned approaches suffer from a serious shortcoming: they require the majority of devices to be honest.
Another approach to achieving solutions with any specified accuracy involves techniques like validation tests (Alistarh et al., 2018; Allen-Zhu et al., 2020) or computation checks (Gorbunov et al., 2022). Nevertheless, they still require a majority of honest devices and rely on strict assumptions in analysis (Alistarh et al., 2018; Allen-Zhu et al., 2020).
While much work on the Byzantines focused on the distributed case with homogeneous data, a different series of papers proposed methods for more realistic training scenarios by allowing data heterogeneity (Wu et al., 2020; El-Mhamdi et al., 2021; Data & Diggavi, 2021; Nguyen et al., 2022). This corresponds, for example, to the federated setting. Methods were primary built around robust aggregation (Karimireddy et al., 2021a; Chang et al., 2019; Data & Diggavi, 2021; Allouah et al., 2024c; Dorfman et al., 2024; Allouah et al., 2024b), and variance reduction techniques (Allouah et al., 2023). One of the most advanced approaches assigned coefficients to clients based on their reliability, using these trust scores to perform gradient steps (Cao et al., 2020; Yan et al., 2024). These studies provided a foundation for Byzantine-robust optimization in the federated setup, but they still suffered from previously mentioned drawbacks.
To circumvent the requirement for a majority of honest workers, several methods attempted to use ground truth data on the server to filter out compromised updates from workers. Thus, Zeno algorithm (Xie et al., 2019), required a validation dataset on the server to evaluate each incoming gradient by comparing it with this dataset. We refer to the function computed on such a validation dataset as the trial function. Nonetheless, Zeno uses this concept only partially. It calculates trust scores using trial functions but then combines the results mostly by regular averaging, except for devices with low trust scores. This results in a significant dependency on the choice of trusted devices. Later, (Cao & Lai, 2019) proposed an alternative algorithm that filters compromised updates. It does this by comparing the updates with a noisy gradient approximation computed on a small dataset, which is effectively equivalent to a validation dataset in Zeno. However, classical averaging techniques, which discard clients with low trust scores, are limited by their high sensitivity to hyperparameters. Specifically, they depend on threshold values for determining malicious devices.
The authors of (Xie et al., 2019; Cao & Lai, 2019) extended their results to handle data heterogeneity by requiring the server to have a representative sample of all device data. However, this assumption is unrealistic in real-world scenarios, undermining the fundamental achievements of federated learning regarding privacy. Moreover, similar approaches (Guo et al., 2021, 2024) also accumulate user data on the server, casting doubt on their applicability.
In addition to these limitations of methods in heterogeneous scenarios, many studies in this field are predominantly heuristic and lack rigorous theoretical analysis (Yan et al., 2024; Guo et al., 2021, 2024; Chang et al., 2019; Xu & Lyu, 2020; Rodríguez-Barroso et al., 2022; Nguyen et al., 2022; Zhang et al., 2022; Huang et al., 2024). In addition, in some studies, the practical component seems to be flawed due to the absence of experiments assessing test accuracy (Gorbunov et al., 2023). Furthermore, theoretical frameworks often do not align with practical part, such as in (Cao et al., 2020), where homogeneous data sampling is assumed despite focusing on the federated learning.
Furthermore, when addressing problems in the distributed or federated setups, local methods (Woodworth et al., 2020; Khaled et al., 2020; Gorbunov et al., 2021b; Nguyen et al., 2022), as well as the partial participation scenario (Yang et al., 2021; Kairouz et al., 2021; Sadiev et al., 2022; Nguyen et al., 2022), is typically assumed. While these options improve computational efficiency and reduce data transmission overhead, only a few studies (Data & Diggavi, 2021; Malinovsky et al., 2023; Allouah et al., 2024a; Dorfman et al., 2024) address these aspects, whereas the majority of works do not. In addition, research is often limited to SGD-like methods, neglecting adaptive algorithms such as Adam (Kingma & Ba, 2014) and RMSProp (Tieleman & Hinton, 2012), which are widely used in machine learning. Such gaps can lead to significant discrepancies between theoretical expectations and practical implementations, impacting the reliability of the results.
Given the challenges and gaps identified in the existing literature, we aim to advance trust scores methodology and trial function concept to enhance the applicability of defense mechanisms.
Contributions. Our main results in tackling Byzantines can be summarized as follows.
$\bullet$ Combine trust scores with the trial function approach. The trial loss function is based on a subset of the training sample stored on the server. To derive trust scores, we evaluate how the gradients sent from devices minimize the trial loss. Two methods embody the ideas of our approach:
  • $(a)$
    The first method assigns weights for the gradients sent from each device based on the extent to which these gradients reduce the trial function in each iteration. In real networks, honest stochastic gradients may increase the target loss. We account for this by incorporating weights from the previous epoch and a momentum parameter for a more stable convergence.
  • $(b)$
    The second approach employs a similar concept of selecting weights. At each step, an additional optimization problem is solved to determine optimal weights that minimize the trial function.
$\bullet$ Milder assumptions. Unlike most existing studies, our approach requires only one reliably honest worker instead of a majority. Moreover, unlike previous trial function-based methods that assume data homogeneity (Cao et al., 2020; Gorbunov et al., 2022), our algorithms work under the more realistic assumption of data similarity in federated learning. In Byzantine literature, it is a common premise for ensuring convergence (Karimireddy et al., 2021a; Gorbunov et al., 2023; Yan et al., 2024).
$\bullet$ Extensions. We adapt our algorithms to practically important scenarios often overlooked in research.
  • $(a)$
    Local methods. In our work, we propose utilizing Local SGD to address the high communication costs typically associated with distributed training.
  • $(b)$
    Partial participation. Our algorithms incorporate the option for partial participation. Thus, devices may not participate in every learning step, and the group involved in the attack can vary across iterations.
  • $(c)$
    Adaptive methods. In this work, we extend our analysis to adaptive algorithms (e.g., Adam and RMSProp), which are widely used in machine learning.
  • $(d)$
    Finding scores from validation. We also introduce a method that employs the concept of a trial function in a slightly different manner. This approach focuses on the trial data stored on the server, where we conduct pairwise validation between the server and the device. Based on the degree of alignment between the predictions made by the device and those generated by the server, we assign weights to each device accordingly.
$\bullet$ Convergence guarantees. We prove upper bounds on convergence rates of main methods and extensions presented for the smooth problem under various assumptions regarding the convexity of the target function (strong convexity, convexity, non-convexity).
$\bullet$ Experiments. We conduct comprehensive experiments, including tests on real data, demonstrating the superiority of our method in both previously studied attacks and scenarios where other methods fail. Our experiments are performed on the CIFAR-10 dataset and real ECG data, using ResNet-18 and ResNet-1D18 neural networks, respectively.
2  Setup
To establish the groundwork for our study, we begin by defining the problem with the assumptions on which our work is based. We consider the problem that is often encountered in distributed machine learning with $\mathcal{D}_{i}$ being an unknown distribution of the training sample data on the $i$-th device:
$\underset{x\in\mathbb{R}^{d}}{\min}\bigl{[}f(x)=\frac{1}{n}\sum\limits_{i=1}^{n}\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{\xi_{i}}(x)\right]\bigr{]}.$ (1)
We consider a setup involving $n$ workers connected to a central server. These workers are divided into two categories: good (or honest) workers, indicated by $\mathcal{G}$ and Byzantines, indicated by $\mathcal{B}$. At each iteration $t$, the sets $\mathcal{G}(t)$ and $\mathcal{B}(t)$ are redefined. This allows the composition of honest and Byzantine workers to vary dynamically over time. At every step, we assume that the set of honest workers $\mathcal{G}(t)$ is nonempty, i.e., $G(t)\overset{def}{=}|\mathcal{G}(t)|\geqslant 1$. The presence of at least one honest worker per iteration maintains the feasibility of the optimization. During training, we do not know the number of Byzantines at each iteration and use it only in theoretical analysis of the worst-case scenario.
To tackle Byzantine attacks, we introduce a pivotal component of our methodology — a trial loss function $\hat{f}$. In the homogeneous setting (1) with $\mathcal{D}_{i}=\mathcal{D}$, we take a separate sample from $\mathcal{D}$ but in a smaller volume than the entire data. The trial function calculated on this data forms $\hat{f}$. Under the heterogeneous data scenario, we sample from the distribution $\mathcal{D}_{1}$ on the server to obtain a delayed data for $\hat{f}$ (indexing the server does not violate generality; we further consider it a device with index 1). Formally, we can write the trial function as $\hat{f}(x)=\frac{1}{N}\sum_{i=1}^{N}f_{1}(x,\xi_{i}),$ where $N$ is the number of samples in $\hat{f}$. This function is stored on the server (obviously an honest device). The sample distribution of the trial function is similar to the entire distribution $\mathcal{D}$ due to the data similarity property. Besides, in practical scenarios, a server may not be able to share the entire dataset, providing only a sample of size $N$. Depending on the size of this sample, $f_{1}$ may differ from $\hat{f}$. Nevertheless, the larger the volume of this delayed sample is, the closer $\hat{f}$ approximates the function $f_{1}$ (discussed in Lemma C.1). Here we outline the assumptions under which we establish the convergence rates.
Assumption 2.1.
The function $\hat{f}$ is $L$-smooth, i.e., $\|\nabla\hat{f}(x)-\nabla\hat{f}(y)\|\leqslant L\|x-y\|$  for any  $x,y\in\mathbb{R}^{d}$.
Assumption 2.2.
The function $\hat{f}$ is:
  1. (a)
    $\mu$-strongly convex if it satisfies the inequality for all $x,y\in\mathbb{R}^{d}$:
    $\hat{f}(y)\geqslant\hat{f}(x)+\left\langle\nabla\hat{f}(x),y-x\right\rangle+\frac{\mu}{2}\|y-x\|^{2}.$
  2. (b)
    convex if it satisfies the inequality for all $x,y\in\mathbb{R}^{d}$:
    $\hat{f}(y)\geqslant\hat{f}(x)+\left\langle\nabla\hat{f}(x),y-x\right\rangle.$
  3. (c)
    non-convex if it has at least one (not necessarily unique) minimum, i.e.,
    $\hat{f}(\hat{x}^{*})=\inf\limits_{x\in\mathbb{R}^{d}}\hat{f}(x)>-\infty.$
Assumption 2.3.
Each worker $i\in\mathcal{G}(t)$ has access to an independent and unbiased stochastic gradient with $\mathbb{E}[g_{i}(x,\xi_{i})]=\nabla f_{i}(x)$ and its variance is bounded by $\sigma^{2}$:
$\mathbb{E}\|g_{i}(x,\xi_{i})-\nabla f_{i}(x)\|^{2}\leqslant\sigma^{2}\quad\text{for all }x\in\mathbb{R}^{d}.$
Assumption 2.4.
We assume data similarity in the following way: good clients possess $(\delta_{1},\delta_{2})$-heterogeneous local loss functions for some $\delta_{1}\geqslant 0$ and $\delta_{2}\geqslant 0$, such that for all $x\in\mathbb{R}^{d}$, the following holds:
$\|\nabla f_{i}(x)-\nabla f(x)\|^{2}\leqslant\delta_{1}+\delta_{2}\|\nabla f(x)\|^{2}~{}~{}~{}~{}\forall i\in\mathcal{G}(t).$
In over-parameterized models, introducing a positive $\delta_{2}$ can sometimes reduce the value of $\delta_{1}$. Several studies on Byzantine-resilient optimization have explored heterogeneous scenarios where honest workers handle distinct local functions (Wu et al., 2020; Karimireddy et al., 2021a). Thus, when heterogeneity is limited to $\delta_{2}$-bounded settings ($\delta_{1}=0$), achieving a predefined accuracy becomes infeasible in the presence of Byzantine workers, even with a trusted Parameter Server (Karimireddy et al., 2021a). In that way, we consider a more general assumption on heterogeneity.
Assumption 2.5.
Byzantine workers are assumed to be omniscient, i.e., they have access to the computations performed by the other workers.
3  Algorithms and Convergence Analysis
3.1  First method: BANT
In this section, we introduce our method designed to counter Byzantine attacks – Algorithm 1, termed Byzantines ANTidote (BANT).
Our method relies on the core idea of assigning trust scores to devices, a technique gaining popularity in defending against attacks (Cao et al., 2020; Yan et al., 2024). We integrate this with the concept of a trial function by aggregating the stochastic gradients $g_{i}^{t}$ of devices with their respective weights $w_{i}^{t}$. To find the latter, we firstly calculate the contribution coefficients for each worker $i$ at each step: ${}\theta_{i}^{t}=\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})~{}$. These coefficients show how the $i$-th device affects the convergence. If $\theta_{i}^{t}>0$, the stochastic gradient minimizes the trial function, and is assigned some weight. Otherwise, it is assigned zero weight (Line 10 in Algorithm 1). We ensure non-negativity with $\left[\theta_{i}^{t}\right]_{0}=\max\{\theta_{i}^{t},0\}$ and normalize to provide a total weight of 1. To address stochastic gradient instability, which can increase the loss function, we introduce a momentum parameter for the weights (Line 10). If all gradients increase the loss, they are given zero weights, stopping the minimization process even without Byzantine devices in the network. By adding momentum, we achieve a more stable convergence in practice. This allows previous good gradients to influence current weights, even if a device receives a small or zero weight in the current iteration. An indicator in the step (Line 14) ensures that gradients maximizing the trial function are ignored, guaranteeing its minimization at each step.
Algorithm 1 BANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$, $\omega_{i}^{0}=\nicefrac{{1}}{{n}}~{}\forall i$
2:  Parameters: Stepsize $\gamma>0$, momentum $\beta\in[0,1]$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Server sends $x^{t}$ to each worker
5:     for all workers $i=1,2,\ldots,n$ in parallel do
6:        Generate $\xi_{i}^{t}$ independently
7:        Compute stochastic gradient $g_{i}^{t}=g_{i}(x^{t},\xi_{i}^{t})$
8:        Send $g_{i}^{t}$ to server
9:     end for
10:     $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})]_{0}}{\sum_{j=1}^{n}[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{j}^{t})]_{0}}$
11:     if each $[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})]_{0}=0$ then
12:        $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{1}{n}$
13:     end if
14:     $x^{t+1}=x^{t}-\gamma\sum_{i=1}^{n}\mathbb{I}_{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})>0]}\omega_{i}^{t}g_{i}^{t}$
15:  end for
16:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
Now we show the convergence results.
Theorem 3.1.
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem (1), after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n}{\beta G}+6L\gamma\sigma^{2}+3\delta_{1}+4\zeta(N).$
The first two terms in the result of Theorem 3.1 replicate the findings from the standard SGD analysis, up to constant factors. The last term, which depends on $\zeta(N)$, is of special interest. The function $\zeta(N)$ reflects the relationship between $f_{1}$ and the trial loss $\hat{f}$ and represents an approximation error. The dependence of $\zeta(N)$ on $N$ is natural: the larger $N$ is, the smaller the error. More precisely, for our function $f$, this error is $\zeta(N)=\mathcal{\widetilde{O}}\left(\frac{1}{N}\right)$ (see Lemma C.1). Although this error degrades convergence, it is actually very common for machine learning tasks. In particular, the original learning problem like (1) is often replaced by its Monte Carlo approximation (Johnson & Zhang, 2013; Defazio et al., 2014; Allen-Zhu, 2018), and the resulting problem is often referred to as the empirical risk minimization (Shalev-Shwartz & Ben-David, 2014). This replacement also leads to an error, but is often not mentioned, considering it a classical and common phenomenon. Finally, $\delta_{1}$ is a typical term which represents data similarity (Assumption 2.4) and is unavoidable in the presence of Byzantines (Wu et al., 2020; Karimireddy et al., 2021a; Gorbunov et al., 2022).
Since our approach bears resemblance with the Zeno method (Xie et al., 2019) in its use of the trial functions, we must mention that we found some issues in their proofs. Specifically, in Theorem 1 of (Xie et al., 2019), the authors incorrectly apply the expectation operator when deriving their recursion. They sample a trial function from the full dataset and use $\mathbb{E}[\hat{f}^{t}(x)]=f(x)$, which is true for a random point. However, in the case of the point $x^{t+1}$, they make an error. Since they sample $\hat{f}^{t}$ in every iteration, the point $x^{t+1}$ depends on the sample $\hat{f}^{t}$, leading to $\mathbb{E}[\hat{f}^{t}(x^{t+1})\mid x^{t}]\neq f(x^{t+1})$. If one attempts to carry through the conditional expectation without considering the full expectation, it becomes impossible to enter the recursion and achieve convergence with respect to the function $f$ itself. To resolve this, we explicitly bound the gradient discrepancy $|\nabla f_{1}(x^{t})-\nabla\hat{f}(x^{t})|$, leveraging the server’s trial function sampling to ensure statistical convergence as sample size grows. This difference in Theorem 3.1 is represented by the discussed $\zeta(N)$. Now we present the final convergence.
Corollary 3.2.
Under the assumptions of Theorem 3.1, for solving the problem (1), after $T$ iterations of Algorithm 1 with $\gamma\leqslant\min\biggl{\{}\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]n}}{\sigma\sqrt{3LG\beta T}}\biggr{\}}$, the following holds:
$\frac{1}{T}\sum\limits_{t=1}^{T-1}$ $\mathbb{E}\|\nabla f(x^{t})\|^{2}=\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}+\frac{\sigma\cdot\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\right).$
If we consider the first two terms in the convergence estimates from Corollary 3.2, the only difference from the classical SGD convergence results Moulines & Bach (2011); Stich (2019) is the additional factor $\frac{n}{G}$, but the rate is asymptotically optimal. The proof of this corollary, the main theorem, as well as the result for the strongly-convex objective, are presented in Appendix, D.
3.2  Second method: AutoBANT
Despite all advantages of the BANT algorithm, it has some imperfections connected to the mechanism of assigning trust scores. The parameter $\beta$ has a negative effect. While it helps honest clients maintain trust scores despite occasional bad gradients, it also allows Byzantine devices to retain their weights during attacks. To combat this, we add an indicator for the trial function reduction (the indication of the device being Byzantine at the considered iteration). However, this limits the theoretical applicability of the method to non-convex problems, prevalent in modern machine learning. To resolve these limitations, we present our second method, called AUxiliary Trial Optimization for Byzantines ANTidote (AutoBANT), formalized as Algorithm 2.
Algorithm 2 AutoBANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$
2:  Parameters: Stepsize $\gamma>0$, error accuracy $\delta$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Server sends $x^{t}$ to each worker
5:     for all workers $i=0,1,2,\ldots,n$ in parallel do
6:        Generate $\xi_{i}^{t}$ independently
7:        Compute stochastic gradient $g_{i}^{t}=g_{i}(x^{t},\xi_{i}^{t})$
8:        Send $g_{i}^{t}$ to server
9:     end for
10:      $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:      $x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
12:  end for
13:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
The idea of assigning weights to devices as a part of the optimization process has recently gained popularity in federated learning. For instance, in many works, it leads to improved solution quality (Li et al., 2023; Tupitsa et al., 2024), or is used in more specific settings such as personalized learning (Mishchenko et al., 2023). We propose to adapt this to Byzantine optimization by optimizing the functionality of $\hat{f}$ with respect to weights calculated after each algorithmic step (Line 10). To solve the minimization problem, we can use various methods, e.g., Mirror Descent (Beck & Teboulle, 2003; Allen-Zhu & Orecchia, 2014):
$\omega^{k+1}~{}=\arg\underset{\omega\in\Delta_{1}^{n}}{\min}$$\Bigl{\{}\eta\left\langle\nabla_{\omega}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{k}g_{i}^{t}\right),\omega\right\rangle+\mathcal{KL}(\omega\|\omega^{k})\Bigr{\}},$
where $\mathcal{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler divergence. The error of solving this is bounded by $\delta$:
$\vspace{-2mm}\left|\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)-\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}\right)\right|\leqslant\delta.$
After solving this auxiliary problem, we produce an actual model update using the optimized weights (Line 11). In light of the proposed method, the question of the cost of implementing such an optimal scheme comes to the forefront. Note that the computational complexity of solving this subproblem at each iteration is only $\mathcal{O}(\nicefrac{{\log n}}{{\delta^{2}}})$ (Beck & Teboulle, 2003), which is not critical.
Theorem 3.3.
Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}<\frac{1}{12}$, 2.5, for solving the problem 1, after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+2\zeta(N)+\frac{4\delta}{\gamma}.$
Corollary 3.4.
Under assumptions of Theorem 3.3, for solving the problem (1), after $T$ iterations of Algorithm 2 with ${\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]G}}{\sigma\sqrt{3LT}}\right\}}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}=\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}{T}+\delta_{1}+\zeta(N)+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}}{\sqrt{TG}}+\left(L+\frac{\sqrt{TL}\sigma}{\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]G}}\right)\delta\right).$
Detailed proofs are presented in Appendix E. Let us discuss the obtained result. In the first and second terms, we see that the method converges the same way as the standard SGD only on honest workers Ghadimi & Lan (2013); Ghadimi et al. (2016). It turns out that we just throw out all Byzantines and this result is almost optimal and unimprovable. As in the case of Algorithm 1, a term responsible for the approximation error $\zeta(N)$ appears in the convergence. By comparing with the result of Corollary 3.2, we improve the rate through a more advanced aggregation mechanism, specifically by removing the factor $\frac{n}{\beta G}$ from the main term and achieving a reduction in variance by a factor of $G$. However, we incurred an additional error $\delta$, which can be seen as a trade-off for solving the subproblem. Furthermore, our approach applies to a broader class of non-convex functions. Addressing the dependence of the stepsize on the number of Byzantine workers, this choice is based on theoretical analysis of the worst-case scenario, considering the number of Byzantine workers. If this number is unknown, setting the minimum possible value of 1 eliminates this dependency.
4  Extensions
Byzantine robust optimization, as discussed earlier, lacks a solid theoretical foundation in several formulations that are the most applicable in the real world. In this work, we address this gap. This section provides a brief overview of the scenarios to which we extend our analysis.
4.1  Local methods
The main idea is that each device performs a predefined number of local steps. Then the aggregation of gradients and mutual updates of model parameters, initialized by the server, take place. This reduces the number of communication rounds. However, it affects the convergence proportionally to the length of the communication round. Complete updates are performed only at specific iterations: $t=t_{k\cdot l}\text{ for some }k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$. During the remaining iterations, we simply perform local updates of the points using the rule $x_{i}^{t+1}=x_{i}^{t}-\gamma g_{i}^{t}$. This approach ensures that communication overhead is minimized while maintaining efficient convergence (see Appendix G).
4.2  Partial participation
It occurs when only a subset of clients actively participates in the training process during each communication round (Yang et al., 2021), allowing clients to join or leave the system. This approach is beneficial in scenarios like mobile edge computing. However, it poses challenges such as incomplete model updates and potential degradation in model performance due to missed contributions from inactive clients (Wang & Ji, 2022; Li et al., 2022). Our methods adapt to the scenario of partial participation due to the assignment of trust scores to devices explicitly participating in training at the considered iteration. Furthermore, it is crucial to account for the minimum number of nodes participating in training across all iterations. Specifically, we analyze $\widetilde{G}(t)=\min_{t\leqslant T}G(t)$, where $G(t)$ denotes the set of active honest workers at iteration $t$. The full version is in Appendix H.
4.3  Scaled methods
Algorithm 3 Scaled AutoBANT (part)
10:  $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:  $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
In modern machine learning, adaptive methods such as Adam (Kingma & Ba, 2014) and RMSProp (Tieleman & Hinton, 2012) became widely popular due to their faster practical convergence and superior performance compared to the standard SGD-like methods. We propose corresponding methods that utilize a diagonal preconditioner $(\hat{P}^{t})^{-1}$, which scales a gradient to $(\hat{P}^{t})^{-1}g_{i}^{t}$, and the step is executed using this scaled gradient. We present the part of our algorithm Scaled AutoBANT, based on Algorithm 2. The estimates we obtain in this analysis are the same as those of the scaled methods in the non-Byzantines regime. A detailed description is in Appendix F.
4.4  Finding scores from validation
Algorithm 4 SimBANT (part)
10:  $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{sim(m(x^{t}-\gamma g_{i}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}))}{\sum\limits_{j=1}^{n}sim(m(x^{t}-\gamma g_{j}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}))}$
11:  $x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
We also propose another method for tackling with Byzantine attacks. The key is that we look not at the loss function $\hat{f}$, as in Algorithms 1, 2, but at the output of the model $m$ on the trial data $\mathcal{\hat{D}}$. The calculation of trust scores $w_{i}$ is now based on how the outputs of the model parameters obtained on the server and on the device are similar to each other. The trust score for $i$-th device is function $\alpha_{i}\rightarrow sim(m(x^{t}-\gamma g_{i}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}))$, which measures the similarity of predictions of server and device models. Based on Algorithm 2, we present the part of the SimBANT algorithm. We provide a detailed explanation in Appendix I.
5  Experiments
To evaluate the performance of the proposed methods, we conduct experiments on two different datasets. The first is the public CIFAR-10 (Krizhevsky et al., 2009), while the second is a proprietary dataset of 12-lead digital electrocardiograms (ECG) from five hospitals. We use ResNet-18 model (He et al., 2016) for CIFAR-10 classification and ResNet1d18 model for ECG abnormality detection. For the CIFAR-10 dataset, we divide the data among ten clients. We consider a homogeneous split with 5,000 images per client, as well as a Dirichlet split with $\alpha=0.5$ and $\alpha=1$. We used 500 samples separately to form $\hat{f}$. For the ECG dataset, we consider five clients, each representing a hospital with 10,000 and 20,000 records. To form $\hat{f}$ on ECG, we use 100 samples from the publicly-available external PTB-XL dataset (Wagner et al., 2020). We solve the task of multiclass classification for CIFAR-10 and binary classification of 4 heart abnormalities for ECG: Atrial FIBrillation (AFIB), First-degree AV Block (1AVB), Premature Ventricular Complex (PVC), and Complete Left Bundle Branch Block (CLBBB). We consider various Byzantine attack scenarios to evaluate our methods.
$\bullet$ Label Flipping. Attackers send gradients based on the loss calculated with randomly flipped labels.
$\bullet$ Sign Flipping. Attackers send the opposite gradient.
$\bullet$ Random Gradients. Attackers send random gradients.
$\bullet$ IPM (Inner Product Manipulation). Attackers send the average gradient of all honest clients multiplied by a factor of -$\kappa$ (we set $\kappa$ to 0.5) (Xie et al., 2020).
$\bullet$ ALIE (A Little Is Enough). Attackers average their gradients and scale the standard deviation to mimic the majority (Baruch et al., 2019).
Algorithm
Without Attack
G-mean f1-score
Label Flipping (60%)
G-mean f1-score
Random Gradients (60%)
G-mean f1-score
IPM (80%)
G-mean f1-score
ALIE (40%)
G-mean f1-score
Adam 0.956$\pm$0.017 0.811$\pm$0.016 0.262$\pm$0.023 0.041$\pm$0.019 0.348$\pm$0.011 0.126$\pm$0.016 0.197$\pm$0.027 0.036$\pm$0.015 0.125$\pm$0.011 0.123$\pm$0.020
FLTrust 0.952$\pm$0.020 0.800$\pm$0.019 0.952$\pm$0.016 0.753$\pm$0.011 0.617$\pm$0.020 0.174$\pm$0.019 0.061$\pm$0.017 0.125$\pm$0.015 0.017$\pm$0.013 0.123$\pm$0.018
Recess 0.949$\pm$0.016 0.783$\pm$0.019 0.366$\pm$0.019 0.128$\pm$0.020 0.593$\pm$0.020 0.163$\pm$0.020 0.493$\pm$0.019 0.112$\pm$0.015 0.450$\pm$0.014 0.127$\pm$0.018
Zeno 0.921$\pm$0.012 0.787$\pm$0.014 0.014$\pm$0.017 0.110$\pm$0.015 0.163$\pm$0.010 0.089$\pm$0.014 0.102$\pm$0.012 0.066$\pm$0.018 0.010$\pm$0.009 0.091$\pm$0.011
CC 0.949$\pm$0.020 0.772$\pm$0.019 0.285$\pm$0.018 0.114$\pm$0.020 0.580$\pm$0.019 0.155$\pm$0.020 0.084$\pm$0.019 0.014$\pm$0.020 0.530$\pm$0.018 0.154$\pm$0.020
CC+fbm 0.954$\pm$0.016 0.808$\pm$0.020 0.840$\pm$0.019 0.716$\pm$0.014 0.562$\pm$0.011 0.151$\pm$0.020 0.027$\pm$0.018 0.123$\pm$0.015 0.876$\pm$0.017 0.594$\pm$0.013
CC+bucketing 0.947$\pm$0.013 0.790$\pm$0.018 0.829$\pm$0.011 0.708$\pm$0.020 0.570$\pm$0.012 0.164$\pm$0.018 0.035$\pm$0.020 0.118$\pm$0.012 0.870$\pm$0.019 0.587$\pm$0.014
Safeguard 0.957$\pm$0.020 0.821$\pm$0.019 0.107$\pm$0.012 0.123$\pm$0.020 0.258$\pm$0.011 0.124$\pm$0.019 0.951$\pm$0.018 0.082$\pm$0.020 0.010$\pm$0.009 0.123$\pm$0.012
VR Marina 0.010$\pm$0.014 0.120$\pm$0.010 0.027$\pm$0.018 0.123$\pm$0.020 0.176$\pm$0.012 0.103$\pm$0.013 0.127$\pm$0.013 0.079$\pm$0.019 0.012$\pm$0.010 0.108$\pm$0.013
BANT 0.953$\pm$0.017 0.830$\pm$0.020 0.956$\pm$0.016 0.777$\pm$0.020 0.948$\pm$0.018 0.809$\pm$0.020 0.946$\pm$0.020 0.676$\pm$0.015 0.947$\pm$0.018 0.770$\pm$0.020
AutoBANT 0.953$\pm$0.019 0.781$\pm$0.020 0.790$\pm$0.020 0.276$\pm$0.020 0.946$\pm$0.019 0.748$\pm$0.018 0.942$\pm$0.020 0.690$\pm$0.020 0.892$\pm$0.016 0.585$\pm$0.020
SimBANT 0.956$\pm$0.020 0.790$\pm$0.018 0.949$\pm$0.020 0.774$\pm$0.020 0.945$\pm$0.020 0.712$\pm$0.018 0.955$\pm$0.020 0.783$\pm$0.018 0.946$\pm$0.019 0.705$\pm$0.020
Table 1: ResNet1d18 on ECG (AFIB) for Byzantine-tolerance techniques under various attacks.
Method Without Attack ALIE (40%) Adam 41.14 $\pm$ 4.20 62.14 $\pm$ 4.83 FLTrust 62.79 $\pm$ 14.53 84.57 $\pm$ 15.17 Recess 86.55 $\pm$ 5.08 98.55 $\pm$ 3.24 Zeno 50.26 $\pm$ 4.35 67.23 $\pm$ 8.42 CC 66.90 $\pm$ 7.65 94.72 $\pm$ 6.17 Safeguard 53.29 $\pm$ 5.11 71.23 $\pm$ 4.67 BANT (ours) 46.17 $\pm$ 3.17 70.21 $\pm$ 5.13 AutoBANT (ours) 62.05 $\pm$ 7.67 83.62 $\pm$ 1.27 SimBANT (ours) 55.47 $\pm$ 6.34 72.49 $\pm$ 7.65
Table 2: Time per communication round for ResNet1d18 on ECG (AFIB)
Figure 1: Test accuracy for ResNet18 on CIFAR-10.
Figure 2: Test accuracy for ResNet18 on Dirichlet.
We define the number of Byzantine clients as a percentage of the total number of clients, and specify it in the attack name. We train BANT and AutoBANT in the scaled version (see Section 4.3) with the Adam preconditioner. We also include SimBANT, Adam, and the existing methods: Zeno (Xie et al., 2019), Recess (Yan et al., 2024), Centered Clip (Karimireddy et al., 2021b), Safeguard (Allen-Zhu et al., 2020), VR Marina (Gorbunov et al., 2021a) and FLTrust (Cao et al., 2020). For Centered Clip, we also added techniques Fixing by Mixing (Allouah et al., 2023) and Bucketing (Karimireddy et al., 2021a). The methods were trained on the CIFAR-10 and ECG datasets for 200 and 150 rounds, respectively. More technical details are presented in Appendix A.1.
The accuracy and loss curves with classification metrics for all considered attacks on the CIFAR-10 test dataset are illustrated in Figures 1, 3 and Table 3 (Appendix A.2), respectively. Figure 1 (see captions (g) and (h)) also shows the accuracy plots of the proposed methods under the Dirichlet distribution with $\alpha=0.5$ for the ALIE and Random Gradients attacks. More details are presented in the Appendix A.3. To assess model performance on the ECG data, we use the G-mean (the square root of sensitivity multiplied by specificity) and the f1-score metrics. Table 1 summarizes the results of the methods for the AFIB disease classification. Detailed results for all considered abnormalities across multiple metrics are presented in Tables 7-10 in Appendix A.2.
Unlike previously established techniques, our methods exhibit robustness against all Byzantine attacks. To further stress test the proposed methods, we consider the most strong Byzantine attacks under statistically heterogeneous setups modeled by the Dirichlet distribution. We also address time per communication round for the methods in Table 2 for the ECG setup. Note that our methods have training times per round comparable to baselines. Round times for the CIFAR-10 setup are presented in Table 4 in Appendix A.2.
We note that AutoBANT performs slightly worse compared to BANT and SimBANT under Random Gradients and ALIE attacks. This occurs due to solving an auxiliary subproblem (Line 10 in Algorithm 2) using Mirror Descent with KL-divergence. According to its properties, the algorithm assigns small but non-zero weights to Byzantines, contributing to unstable convergence, while BANT and SimBANT lack this drawback.
Recess and FLTrust leverage the concept of trust scores but rely on the majority of honest devices. As a result, it leads to a significant decrease in final quality under the majority of Byzantines in Random Gradients and IPM setups, as well as with ALIE attack that simulates a malicious majority. Similar behavior is observed for the CC and Safeguard methods, which also suffer from sensitivity to parameter tuning. Fixing by Mixing and Bucketing increase Label Flipping and ALIE metrics for ECG setup, but do not provide reliable convergence for all cases. Zeno exploits the trial function approach, but it relies on the number of Byzantine clients. We address the choice of this hyperparameter in Table 5 (Appendix A.2). In contrast, BANT-based methods demonstrate resilience to extreme attacks while enabling adaptive selection of trusted clients.
Acknowledgements
A significant part of this work was carried out while G. Molodtsov and D. Medyakov were visiting research assistants at Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
Appendix
Appendix A Additional Experiments
A.1  Technical details
Our implementation is developed in Python 3.10. We simulate a distributed system on a single server. The server is equipped with an AMD EPYC 7513 32-Core Processor running at 2.6 GHz and Nvidia A100 SXM4 40GB. This configuration is used for the experiments described in Section 5.
To ensure a fair comparison, we maintain consistent hyperparameters across all methods. The batch size is set to 32 for CIFAR-10 and 64 for the ECG dataset, with a local client learning rate of 0.003 and adam preconditioner. All clients perform local computations for 1 epoch. For CIFAR-10, cross-entropy was used as the local client loss function, while for ECG is was its binary version. To address the problem of imbalance of positive and negative examples in the ECG, the minority class was reweighted to the majority in the corresponding loss function. For the ECG classification task, we train the models on 10-second 12-lead ECG records, with all records resampled to a frequency of 500 Hz. Additionally, we train the models exclusively on patients older than 18 years of age.
We also select specific parameters for the implemented methods. For the Safeguard method, we use window sizes of 1 and 6 for two different accumulation settings, with the threshold chosen automatically, as described in the original paper. We adapt the CC method to a local computation case and set the clipping coefficient $\tau=0.1$ and the SGD momentum $\beta=0.9$ for CIFAR-10, as well as $\tau=1$ and $\beta=0.5$ for the ECG case. For both setups we fixed the number of clipping iterations to $l=1$. For the Fixing-by-Mixing technique, we set the number of Byzantine clients $f$ to be less than half of all clients as suggested in the article: for the CIFAR-10 $f=4$, while for the ECG $f=2$. For the Bucketing technique we apply 2-bucketing strategy with global learning rate $\eta=0.9$. For the Recess method, we set the decrease score equal to $0.1$. For the Zeno method, we set the regularization weight $\rho=0.0005$ as a default value in the paper and the trim parameter $b=2$ for the ECG setup and $b=3$ for CIFAR-10. We address the choice of hyperparameter $b$ in Table 5 in Appendix A.2, as it is critical in real-world scenarios. For the BANT method, we set the momentum parameter $\beta=0.5$. The AutoBANT method uses the number of optimization epochs equal to 5 and $\gamma=1$. For SimBANT, we set the softmax temperature parameter for the model logits to 0.05.
Additionally, we define distinct functions for SimBANT based on the dataset, as ECG classification is binary, whereas CIFAR-10 is a multi-class classification problem. Specifically, for the ECG dataset, we use:
$sim_{\text{ECG}}(x,y)=1-|x-y|,$ (2)
where $x$ is the output of the client model, and $y$ is the output of the model fine-tuned on the server. For the CIFAR-10 dataset, we apply cosine similarity:
$sim_{\text{CIFAR}}(x,y)=\frac{x\cdot y}{\|x\|\|y\|},$ (3)
where $y$ is one-hot encode targets.
A.2  CIFAR-10 Experiments
In this section, we present supplementary data regarding the experiments conducted. As mentioned in the main part of the paper, we utilized ResNet-18 models on the CIFAR-10 dataset. We begin with a comparative Table 3 of the methods applied to the CIFAR-10 dataset.
Algorithm Without Attack Label Flipping (50%) Sign Flipping (60%) Random Gradients (50%) IPM (70%) ALIE (40%) Sign Flipping (40%) IPM (50%) Adam 0.902 0.207 0.100 0.100 0.100 0.100 0.624 0.832 FLTrust 0.767 0.694 0.100 0.100 0.100 0.100 0.254 0.519 Recess 0.887 0.633 0.100 0.103 0.106 0.128 0.488 0.774 Zeno 0.910 0.156 0.410 0.100 0.100 0.100 0.838 0.100 CC 0.917 0.603 0.102 0.100 0.100 0.100 0.511 0.864 CC+fbm 0.915 0.887 0.098 0.100 0.101 0.100 0.823 0.923 CC+bucketing 0.845 0.818 0.089 0.101 0.100 0.101 0.815 0.100 Safeguard 0.918 0.102 0.100 0.102 0.104 0.113 0.826 0.112 VR Marina 0.100 0.100 0.100 0.100 0.100 0.100 0.100 0.100 BANT 0.864 0.861 0.846 0.846 0.725 0.856 0.856 0.751 AutoBANT 0.906 0.884 0.783 0.898 0.666 0.882 0.839 0.847 SimBANT 0.909 0.855 0.827 0.865 0.623 0.878 0.852 0.827
Table 3: ResNet18 on CIFAR-10. Accuracy under various attacks.
This table illustrates the performance of different algorithms under various attacks, highlighting their effectiveness. We’ve supplemented it with scenarios of Sign Flipping ($40\%$) and IPM ($50\%$) attacks compared to the main part to demonstrate the results of baselines in less stressfull conditions. It is noteworthy that while existing methods provide some level of protection against certain attacks, none are effective when faced with a majority of malicious clients in gradient attacks, as well as in IPM ($70\%$) and ALIE ($40\%$), which simulates the majority. In contrast, all three of our methods demonstrate impressive performance in such attack scenarios. Furthermore, it is important to compare these results with those in the first column, which represents the metric in the absence of attacks. As we can see, our methods achieve only a slight reduction in the metric, yet they maintain relatively strong performance even under the most severe attacks. This resilience underscores the effectiveness of our approaches and their potential for real-world applications where robust defense mechanisms are crucial.
Now we examine the graphs depicting the reduction of loss over the course of training (Figure 3). The result is similar: as the number of attackers increases, the existing methods exhibit divergence, while our methods continue to decrease the loss effectively. As mentioned in the main part of the paper, AutoBANT may behave inconsistently under Random Gradients and ALIE attacks. This instability can be attributed to the solution of an additional minimization problem and the absence of an indicator that prevents theoretical advancement in non-convex scenarios. Nevertheless, even under these circumstances, AutoBANT demonstrates significantly better results compared to its counterparts.
Figure 3: ResNet18 on CIFAR-10. Loss on test for Byzantine-tolerance techniques under various attacks.
Algorithm Without Attack ALIE (40%) Adam 17.76 ± 1.76 30.74 ± 1.87 FLTrust 57.64 ± 7.55 72.75 ± 1.82 Recess 65.88 ± 16.02 76.99 ± 6.27 Zeno 38.13 ± 2.26 54.88 ± 4.71 CC 22.95 ± 2.30 43.34 ± 6.28 CC+fbm 23.44 ± 2.81 47.16 ± 7.95 CC+bucketing 20.18 ± 2.12 43.46 ± 6.19 Safeguard 43.26 ± 3.14 58.29 ± 2.31 VR Marina 55.84 ± 9.14 73.91 ± 6.89 BANT 29.31 ± 2.56 55.89 ± 6.29 AutoBANT 49.25 ± 5.77 67.27 ± 3.49 SimBANT 33.40 ± 5.00 60.11 ± 8.70
Table 4: Time (in seconds) per communication round for ResNet18 on CIFAR-10 without attack and under ALIE attack.
Table 4 reports the average time per communication round, including standard deviation for various federated learning algorithms using ResNet18 on CIFAR-10 in two settings: Without attack and ALIE. We see that Recess took the longest time, since it requires double local calculations of the client in one round of communication. AutoBANT requires a little more time compared to baselines, except for the Recess and FLTrust methods. We attribute this to the solution of the auxiliary task Algorithm 4.3, Line 10), which is slightly longer than in the ECG setup and has a greater impact on the execution time. However, it is not expected that the size of $\hat{f}$ increases significantly as the system scales. Thus, this contribution is negligible in real-world scenarios, as demonstrated by the ECG experiment in Table 2. In addition, these computations are performed on the central node, which has more computational resources in the federated learning paradigm. All of this reflects the applicability of the proposed methods in real-world setups.
A.3  Dirichlet Distributions
To examine the robustness of the proposed methods under a heterogeneous setup, we perform Byzantine attacks on CIFAR-10 with Dirichlet distribution of client data. So $\mathbf{p}_{k}\sim\text{Dir}(\alpha\cdot\mathbf{1}_{n})$ represents distribution of samples with label $k=\overline{1,K}$ over $n$ clients, i.e. $\mathbf{p}_{k}\in S_{n}(1)$ where $S_{n}(1)$ is the standard $n$ unit simplex. The hyperparameter $\alpha$ is called the concentration parameter, and it controls the degree of heterogeneity of the client data distribution. We address two heterogeneity regimes: medium with $\alpha=1$ and strong with $\alpha=0.5$. Since the goal of the experiment is to test the robustness of the proposed methods, we consider the 3 strongest attacks in our setup: Random Gradient (50%), ALIE (40%), and IPM (70%). Figure 4, 5 illustrates the performance of the BANT-based methods.
Figure 4: Test accuracy for ResNet18 on CIFAR-10 with Dirichlet heterogeneity.
Figure 5: Test loss for ResNet18 on CIFAR-10 with Dirichlet heterogeneity.
A.4  ECG Experiments
Now, let us move on to the choice of the trim hyperparameter $b$ in the Zeno method. This parameter cuts off the number of clients that will not be used in the aggregation of the global model in the training round. The ranking of clients is based on $Score_{\gamma,\rho}$, which exploits the trial function approach. In the paper, it is assumed that $b\geqslant q$, where $q$ is the number of Byzantines participating in the training. Thus, the method provides protection against an arbitrary number of malicious participants and can be stable in critical attack setups. However, in real-world scenarios, we do not have a priori information about the number of Byzantines. As a result, Zeno does not show comparable quality metrics in the ECG setup for $b=2$ compared to the BANT-like methods, which is reflected in Table 1.
Algorithm
Without Attack
G-mean f1-score
Label Flipping (60%)
G-mean f1-score
Random Gradients (60%)
G-mean f1-score
IPM (80 %)
G-mean f1-score
ALIE (40 %)
G-mean f1-score
Zeno ($b$ = 3) 0.953$\pm$0.018 0.806$\pm$0.017 0.950$\pm$0.020 0.753$\pm$0.019 0.954$\pm$0.020 0.770$\pm$0.017 0.945$\pm$0.017 0.730$\pm$0.020 0.946$\pm$0.015 0.717$\pm$0.016
Zeno 0.921$\pm$0.012 0.787$\pm$0.014 0.014$\pm$0.017 0.110$\pm$0.015 0.163$\pm$0.010 0.089$\pm$0.014 0.102$\pm$0.012 0.066$\pm$0.018 0.010$\pm$0.009 0.091$\pm$0.011
Table 5: ResNet1d18 on ECG (AFIB). G-mean and f1-score for Zeno with different hyperparameters under attacks.
We address the issue of choosing this hyperparameter in the ECG setup in Table 5. As can be seen from the results, a suitably selected trim coefficient value of $b=3$, which cuts off $60\%$ of Byzantine clients, shows comparable results with the BANT-like methods in all considered attack scenarios.
Algorithm
Sign Flipping (60 %)
G-mean f1-score
IPM (60%)
G-mean f1-score
Adam 0.304$\pm$0.015 0.116$\pm$0.018 0.952$\pm$0.014 0.738$\pm$0.011
FLTrust 0.586$\pm$0.018 0.179$\pm$0.015 0.011$\pm$0.014 0.123$\pm$0.013
Recess 0.359$\pm$0.018 0.115$\pm$0.014 0.933$\pm$0.017 0.611$\pm$0.018
Zeno 0.017$\pm$0.016 0.130$\pm$0.018 0.010$\pm$0.018 0.140$\pm$0.020
CC 0.479$\pm$0.020 0.124$\pm$0.017 0.945$\pm$0.020 0.710$\pm$0.016
CC+fbm 0.155$\pm$0.016 0.124$\pm$0.017 0.948$\pm$0.018 0.695$\pm$0.020
CC+bucketing 0.137$\pm$0.017 0.119$\pm$0.010 0.944$\pm$0.016 0.689$\pm$0.013
Safeguard 0.084$\pm$0.016 0.014$\pm$0.018 0.109$\pm$0.010 0.123$\pm$0.016
VR Marina 0.096$\pm$0.017 0.078$\pm$0.019 0.098$\pm$0.011 0.110$\pm$0.014
BANT 0.943$\pm$0.019 0.792$\pm$0.019 0.949$\pm$0.020 0.704$\pm$0.017
AutoBANT 0.737$\pm$0.020 0.243$\pm$0.019 0.948$\pm$0.018 0.695$\pm$0.015
SimBANT 0.951$\pm$0.020 0.760$\pm$0.020 0.965$\pm$0.017 0.753$\pm$0.020
Table 6: ResNet1d18 on ECG (AFIB). G-mean and f1-score for Byzantine-tolerance techniques under 2 attacks.
Here, we will present all the obtained metric results for the four heart abnormalities we investigated: Atrial FIBrillation (AFIB), First-degree AV block (1AVB), Premature Ventricular Complex (PVC), and Complete Left Bundle Branch Block (CLBBB). In Table 6, we present results for AFIB pathology under 2 additional attacks scenarios: Sign Flipping $(60\%)$ and IPM $(60\%)$. In addition, we compare all methods for the attacks described in 5 section in Tables 7 - 10. To obtain confidence intervals in Tables 1, 5, 6, each run was repeated $5$ times. We did not estimate the variance of metrics for the 1AVB, PVC and CLBBB pathologies in Tables 8-10 due to computational challenges.
Algorithm Sensitivity Specificity G-mean f1-score
Without Attack Adam 0.940 0.974 0.956 0.811
FLTrust 0.932 0.972 0.952 0.800
Recess 0.929 0.969 0.949 0.783
Zeno 0.890 0.952 0.921 0.787
CC 0.932 0.966 0.949 0.772
CC+fbm 0.930 0.978 0.954 0.808
CC+bucketing 0.910 0.985 0.947 0.790
Safeguard 0.940 0.976 0.957 0.821
VR Marina 0.150 0.001 0.010 0.120
BANT 0.929 0.978 0.953 0.830
AutoBANT 0.940 0.967 0.953 0.781
SimBANT 0.943 0.969 0.956 0.790
Label Flip (60%) Adam 0.089 0.774 0.262 0.041
FLTrust 0.943 0.961 0.952 0.753
Recess 0.146 0.920 0.366 0.128
Zeno 0.160 0.001 0.014 0.110
CC 0.840 0.097 0.285 0.114
CC+fbm 0.880 0.801 0.840 0.716
CC+bucketing 0.870 0.789 0.829 0.708
Safeguard 0.018 0.063 0.107 0.123
VR Marina 0.210 0.003 0.027 0.123
BANT 0.947 0.966 0.956 0.777
AutoBANT 0.964 0.647 0.790 0.276
SimBANT 0.932 0.967 0.949 0.774
Sign Flip (60%) Adam 0.096 0.961 0.304 0.116
FLTrust 0.466 0.738 0.586 0.179
Recess 0.142 0.906 0.359 0.115
Zeno 0.230 0.001 0.017 0.130
CC 0.328 0.701 0.479 0.124
CC+fbm 0.270 0.089 0.155 0.124
CC+bucketing 0.250 0.075 0.137 0.119
Safeguard 0.007 0.998 0.084 0.014
VR Marina 0.150 0.061 0.096 0.078
BANT 0.915 0.972 0.943 0.792
AutoBANT 0.854 0.636 0.737 0.243
SimBANT 0.940 0.963 0.951 0.760
Random Gradients (60%) Adam 0.136 0.893 0.348 0.126
FLTrust 0.623 0.611 0.617 0.174
Recess 0.562 0.626 0.593 0.163
Zeno 0.300 0.088 0.163 0.089
CC 0.459 0.733 0.580 0.155
CC+fbm 0.500 0.632 0.562 0.252
CC+bucketing 0.420 0.774 0.570 0.164
Safeguard 0.929 0.071 0.258 0.124
VR Marina 0.320 0.096 0.176 0.103
BANT 0.922 0.975 0.948 0.809
AutoBANT 0.932 0.961 0.946 0.748
SimBANT 0.951 0.940 0.945 0.712
IPM (60 %) Adam 0.947 0.957 0.952 0.738
FLTrust 0.210 0.001 0.011 0.123
Recess 0.947 0.919 0.933 0.611
Zeno 0.200 0.001 0.010 0.140
CC 0.940 0.950 0.945 0.710
CC+fbm 0.950 0.946 0.948 0.695
CC+bucketing 0.940 0.948 0.944 0.689
Safeguard 0.190 0.062 0.109 0.123
VR Marina 0.170 0.056 0.098 0.110
BANT 0.950 0.947 0.949 0.704
AutoBANT 0.950 0.945 0.948 0.695
SimBANT 0.974 0.955 0.965 0.753
IPM (80 %) Adam 0.043 0.904 0.197 0.036
FLTrust 0.220 0.016 0.061 0.124
Recess 0.456 0.533 0.493 0.112
Zeno 0.131 0.080 0.102 0.066
CC 0.007 0.997 0.084 0.014
CC+fbm 0.210 0.003 0.027 0.123
CC+bucketing 0.220 0.005 0.035 0.118
Safeguard 0.200 0.060 0.110 0.082
VR Marina 0.240 0.067 0.127 0.079
BANT 0.954 0.939 0.946 0.676
AutoBANT 0.940 0.945 0.942 0.690
SimBANT 0.950 0.960 0.955 0.783
ALIE (40%) Adam 0.265 0.058 0.125 0.123
FLTrust 0.220 0.001 0.017 0.123
Recess 0.249 0.811 0.450 0.127
Zeno 0.180 0.001 0.010 0.091
CC 0.370 0.758 0.530 0.154
CC+fbm 0.870 0.882 0.876 0.594
CC+bucketing 0.890 0.850 0.870 0.587
Safeguard 0.150 0.000 0.010 0.123
VR Marina 0.200 0.001 0.012 0.108
BANT 0.929 0.966 0.947 0.770
AutoBANT 0.861 0.924 0.892 0.585
SimBANT 0.943 0.949 0.946 0.705
Table 7: ResNet1d18 on ECG (AFIB).
Algorithm Sensitivity Specificity G-mean f1-score
Without Attack Adam 0.896 0.871 0.884 0.335
FLTrust 0.857 0.883 0.870 0.344
Recess 0.909 0.870 0.889 0.337
Zeno 0.890 0.883 0.886 0.353
CC 0.864 0.896 0.880 0.371
CC+fbm 0.888 0.868 0.877 0.353
CC+bucketing 0.869 0.887 0.877 0.351
Safeguard 0.877 0.883 0.880 0.350
VR Marina 0.825 0.922 0.872 0.420
BANT 0.831 0.916 0.873 0.408
AutoBANT 0.825 0.922 0.872 0.420
FineTuned 0.851 0.894 0.872 0.362
Label Flip (60%) Adam 0.210 0.002 0.022 0.069
FLTrust 0.870 0.903 0.886 0.388
Recess 0.210 0.001 0.017 0.113
Zeno 0.400 0.039 0.125 0.309
CC 0.290 0.120 0.187 0.112
CC+fbm 0.725 0.866 0.792 0.265
CC+bucketing 0.740 0.818 0.778 0.305
Safeguard 0.100 0.001 0.010 0.069
VR Marina 0.220 0.002 0.023 0.125
BANT 0.812 0.934 0.871 0.455
AutoBANT 0.877 0.895 0.886 0.373
FineTuned 0.890 0.857 0.873 0.311
Sign Flip (60%) Adam 0.597 0.685 0.640 0.119
FLTrust 0.007 0.980 0.080 0.008
Recess 0.058 0.874 0.226 0.026
Zeno 0.190 0.079 0.123 0.015
CC 1.000 0.002 0.041 0.070
CC+fbm 0.260 0.080 0.145 0.122
CC+bucketing 0.250 0.076 0.138 0.117
Safeguard 0.981 0.030 0.170 0.070
VR Marina 0.130 0.058 0.087 0.065
BANT 0.669 0.951 0.797 0.447
AutoBANT 0.916 0.845 0.879 0.301
FineTuned 0.617 0.616 0.616 0.103
Random Gradients (60%) Adam 0.220 0.005 0.036 0.074
FLTrust 1.000 0.001 0.027 0.069
Recess 0.468 0.757 0.595 0.117
Zeno 0.220 0.061 0.116 0.098
CC 0.909 0.227 0.455 0.080
CC+fbm 0.610 0.390 0.487 0.362
CC+bucketing 0.650 0.416 0.520 0.340
Safeguard 1.000 0.020 0.140 0.071
VR Marina 0.230 0.057 0.115 0.103
BANT 0.805 0.919 0.860 0.405
AutoBANT 0.896 0.864 0.880 0.324
FineTuned 0.857 0.890 0.873 0.357
IPM (60 %) Adam 0.933 0.622 0.762 0.154
FLTrust 0.013 0.964 0.112 0.013
Recess 0.922 0.852 0.886 0.313
Zeno 0.260 0.169 0.210 0.098
CC 0.571 0.649 0.609 0.104
CC+fbm 0.700 0.480 0.580 0.130
CC+bucketing 0.610 0.374 0.478 0.142
Safeguard 0.050 0.001 0.008 0.078
VR Marina 0.200 0.036 0.085 0.115
BANT 0.721 0.945 0.825 0.449
AutoBANT 0.890 0.865 0.877 0.323
FineTuned 0.857 0.893 0.875 0.363
IPM (80 %) Adam 0.312 0.580 0.425 0.050
FLTrust 0.201 0.751 0.389 0.051
Recess 0.558 0.508 0.533 0.076
Zeno 0.240 0.065 0.125 0.110
CC 0.857 0.310 0.515 0.084
CC+fbm 0.220 0.220 0.038 0.113
CC+bucketing 0.230 0.008 0.045 0.107
Safeguard 0.481 0.499 0.490 0.064
VR Marina 0.260 0.063 0.128 0.083
BANT 0.857 0.884 0.870 0.345
AutoBANT 0.851 0.900 0.875 0.375
FineTuned 0.903 0.876 0.889 0.344
ALIE (40%) Adam 1.000 0.001 0.031 0.069
FLTrust 0.220 0.032 0.085 0.150
Recess 0.220 0.055 0.110 0.065
Zeno 0.220 0.025 0.075 0.124
CC 0.520 0.443 0.480 0.115
CC+fbm 0.690 0.626 0.657 0.350
CC+bucketing 0.700 0.585 0.640 0.380
Safeguard 1.000 0.000 0.000 0.069
VR Marina 0.200 0.001 0.010 0.103
BANT 0.935 0.850 0.892 0.314
AutoBANT 0.818 0.888 0.852 0.339
SimBANT 0.968 0.818 0.890 0.282
Table 8: ResNet1d18 on ECG (1AVB).
Algorithm Sensitivity Specificity G-mean f1-score
Without Attack Adam 0.977 0.974 0.975 0.790
FLTrust 0.972 0.971 0.972 0.772
Recess 0.972 0.961 0.967 0.720
Zeno 0.977 0.975 0.976 0.801
CC 0.972 0.959 0.965 0.707
CC+fbm 0.960 0.980 0.970 0.755
CC+bucketing 0.950 0.986 0.968 0.740
Safeguard 0.977 0.965 0.971 0.743
VR Marina 0.240 0.060 0.120 0.098
BANT 0.931 0.981 0.955 0.810
AutoBANT 0.972 0.970 0.971 0.765
FineTuned 0.963 0.971 0.967 0.770
Label Flip (60%) Adam 0.639 0.710 0.673 0.180
FLTrust 0.220 0.059 0.114 0.096
Recess 0.931 0.937 0.934 0.596
Zeno 0.310 0.037 0.108 0.210
CC 0.005 0.999 0.068 0.009
CC+fbm 0.930 0.920 0.925 0.760
CC+bucketing 0.910 0.910 0.910 0.735
Safeguard 0.981 0.971 0.976 0.782
VR Marina 0.200 0.036 0.085 0.115
BANT 0.870 0.979 0.923 0.767
AutoBANT 0.972 0.976 0.974 0.805
FineTuned 0.944 0.957 0.951 0.686
Sign Flip (60%) Adam 0.639 0.710 0.673 0.180
FLTrust 0.220 0.032 0.085 0.118
Recess 0.931 0.937 0.934 0.596
Zeno 0.190 0.069 0.115 0.010
CC 0.005 0.999 0.068 0.009
CC+fbm 0.250 0.072 0.135 0.117
CC+bucketing 0.230 0.055 0.113 0.096
Safeguard 0.981 0.971 0.976 0.782
VR Marina 0.120 0.035 0.065 0.010
BANT 0.870 0.979 0.923 0.767
AutoBANT 0.972 0.976 0.974 0.805
FineTuned 0.944 0.957 0.951 0.686
Random Gradients (60%) Adam 0.220 0.005 0.035 0.120
FLTrust 0.220 0.063 0.118 0.110
Recess 0.255 0.845 0.464 0.122
Zeno 0.180 0.053 0.098 0.010
CC 0.250 0.048 0.110 0.130
CC+fbm 0.270 0.083 0.150 0.117
CC+bucketing 0.220 0.043 0.098 0.123
Safeguard 0.009 0.997 0.096 0.017
VR Marina 0.120 0.037 0.067 0.010
BANT 0.944 0.969 0.957 0.747
AutoBANT 0.963 0.978 0.970 0.809
FineTuned 0.963 0.945 0.954 0.644
IPM (60 %) Adam 0.796 0.976 0.882 0.708
FLTrust 0.240 0.046 0.106 0.114
Recess 0.944 0.938 0.941 0.608
Zeno 0.220 0.025 0.075 0.124
CC 0.944 0.933 0.939 0.590
CC+fbm 0.930 0.960 0.945 0.625
CC+bucketing 0.930 0.950 0.940 0.610
Safeguard 0.782 0.217 0.412 0.095
VR Marina 0.190 0.037 0.084 0.115
BANT 0.931 0.977 0.954 0.790
AutoBANT 0.949 0.971 0.960 0.762
FineTuned 0.954 0.965 0.959 0.729
IPM (80 %) Adam 0.958 0.005 0.069 0.093
FLTrust 0.005 0.991 0.068 0.008
Recess 0.426 0.502 0.462 0.079
Zeno 0.300 0.225 0.260 0.108
CC 0.463 0.265 0.350 0.061
CC+fbm 0.220 0.010 0.047 0.114
CC+bucketing 0.200 0.006 0.035 0.105
Safeguard 0.065 0.831 0.232 0.031
VR Marina 0.210 0.063 0.115 0.086
BANT 0.931 0.980 0.955 0.807
AutoBANT 0.970 0.965 0.968 0.820
FineTuned 0.963 0.937 0.950 0.611
ALIE (40%) Adam 0.200 0.021 0.065 0.120
FLTrust 0.398 0.581 0.481 0.086
Recess 0.200 0.028 0.075 0.135
Zeno 0.190 0.063 0.110 0.078
CC 0.913 0.116 0.325 0.084
CC+fbm 0.880 0.821 0.850 0.580
CC+bucketing 0.890 0.870 0.880 0.610
Safeguard 0.218 0.881 0.438 0.126
VR Marina 0.240 0.064 0.124 0.105
BANT 0.954 0.955 0.954 0.680
AutoBANT 0.954 0.981 0.967 0.826
FineTuned 0.917 0.954 0.935 0.658
Table 9: ResNet1d18 on ECG (PVC).
Algorithm Sensitivity Specificity G-mean f1-score
Without Attack Adam 0.979 0.957 0.968 0.508
FLTrust 0.990 0.947 0.968 0.462
Recess 0.969 0.963 0.966 0.538
Zeno 0.990 0.957 0.973 0.509
CC 0.979 0.945 0.962 0.445
CC+fbm 0.970 0.970 0.970 0.480
CC+bucketing 0.950 0.950 0.950 0.465
Safeguard 0.990 0.961 0.975 0.537
VR Marina 0.220 0.021 0.068 0.117
BANT 0.990 0.947 0.968 0.460
AutoBANT 0.989 0.936 0.962 0.415
FineTuned 0.990 0.953 0.971 0.491
Label Flip (60%) Adam 0.220 0.065 0.120 0.115
FLTrust 0.198 0.630 0.353 0.023
Recess 0.021 0.873 0.135 0.006
Zeno 0.190 0.084 0.127 0.005
CC 0.764 0.854 0.808 0.199
CC+fbm 0.860 0.800 0.830 0.210
CC+bucketing 0.790 0.830 0.810 0.214
Safeguard 0.989 0.958 0.974 0.517
VR Marina 0.260 0.042 0.105 0.125
BANT 0.990 0.923 0.956 0.370
AutoBANT 0.990 0.955 0.972 0.502
FineTuned 0.990 0.952 0.971 0.487
Sign Flip (60%) Adam 1.000 0.001 0.027 0.044
FLTrust 0.865 0.747 0.804 0.134
Recess 0.220 0.102 0.150 0.044
Zeno 0.250 0.193 0.220 0.054
CC 0.260 0.177 0.215 0.117
CC+fbm 0.320 0.204 0.256 0.117
CC+bucketing 0.300 0.052 0.125 0.240
Safeguard 0.698 0.617 0.656 0.076
VR Marina 0.210 0.029 0.078 0.112
BANT 0.979 0.944 0.962 0.444
AutoBANT 0.969 0.934 0.951 0.401
FineTuned 0.969 0.958 0.964 0.512
Random Gradients (60%) Adam 1.000 0.007 0.085 0.044
FLTrust 1.000 0.003 0.054 0.044
Recess 0.427 0.963 0.641 0.282
Zeno 0.310 0.054 0.130 0.210
CC 0.130 0.120 0.125 0.044
CC+fbm 0.310 0.201 0.250 0.130
CC+bucketing 0.190 0.170 0.180 0.054
Safeguard 0.052 0.987 0.227 0.064
VR Marina 0.210 0.030 0.080 0.112
BANT 0.990 0.948 0.969 0.465
AutoBANT 0.990 0.931 0.960 0.396
FineTuned 0.990 0.971 0.980 0.607
IPM (60 %) Adam 0.979 0.957 0.968 0.511
FLTrust 0.979 0.296 0.538 0.060
Recess 0.958 0.945 0.951 0.438
Zeno 0.330 0.027 0.095 0.215
CC 0.958 0.950 0.954 0.465
CC+fbm 0.950 0.940 0.945 0.440
CC+bucketing 0.940 0.936 0.938 0.460
Safeguard 0.979 0.539 0.727 0.089
VR Marina 0.230 0.041 0.098 0.115
BANT 0.813 0.982 0.893 0.629
AutoBANT 0.979 0.941 0.960 0.429
FineTuned 0.938 0.930 0.934 0.376
IPM (80 %) Adam 0.073 0.506 0.192 0.006
FLTrust 0.330 0.017 0.075 0.210
Recess 0.042 0.952 0.199 0.026
Zeno 0.230 0.141 0.180 0.048
CC 0.073 0.884 0.254 0.024
CC+fbm 0.190 0.075 0.120 0.065
CC+bucketing 0.220 0.200 0.210 0.020
Safeguard 1.000 0.002 0.046 0.044
VR Marina 0.220 0.005 0.035 0.118
BANT 0.990 0.951 0.970 0.481
AutoBANT 0.979 0.947 0.963 0.459
FineTuned 0.990 0.943 0.966 0.441
ALIE (40%) Adam 0.250 0.176 0.210 0.123
FLTrust 0.220 0.089 0.140 0.065
Recess 1.000 0.005 0.071 0.044
Zeno 0.220 0.032 0.085 0.114
CC 0.460 0.599 0.525 0.160
CC+fbm 0.490 0.470 0.480 0.210
CC+bucketing 0.540 0.463 0.500 0.120
Safeguard 0.220 0.060 0.115 0.087
VR Marina 0.200 0.078 0.125 0.072
BANT 0.979 0.968 0.973 0.578
AutoBANT 1.000 0.814 0.902 0.198
FineTuned 0.958 0.957 0.958 0.501
Table 10: ResNet1d18 on ECG (CLBBB).
Appendix B Notation
The following sections will be dedicated to the theoretical proof of all aspects discussed in the main body of the paper. A complete table of contents can be found on page 13.
In order to facilitate the understanding of the proofs presented in the appendix, as well as to simplify the interaction with all the formulas throughout this paper, we provide a comprehensive list of notation used in this study in the form of the following table.
$N$ Number of samples in the trial function $\gamma$ Learning rate (stepsize) in optimization $\hat{f}$ Trial loss function on the server $\mathbb{I}_{[a>0]}$ Indicator function taking value 1 if $a>0$, otherwise 0
$d$ Dimensionality of the parameter space $\mathbb{R}^{d}$ $L$ Smoothness constant (2.1) $f$ Objective function $f(x)$ in distributed learning $\hat{P}^{t}$ Adaptive preconditioner matrix at iteration $t$
$n=n(t)$ Number of workers in the distributed system $\mu$ Strong convexity constant (2.2(a)) $f_{i}$ Local objective function on $i$-th worker $\langle a,b\rangle_{\hat{P}}$ Weighted inner product with $\hat{P}$
$\mathcal{G}=\mathcal{G}(t)$ Set of honest workers $\beta$ Momentum parameter for weights $\omega_{i}^{t}$ $f_{1}$ Local objective function on the server $\alpha$ Lower bound on the preconditioner (F.1)
$G=G(t)$ Number of honest workers $\delta$ Approximation error in $\arg\min$ finding (2) $g_{i}$ Stochastic gradient of worker $i$ $\Gamma$ Upper bound on the preconditioner (F.1)
$\mathcal{B}=\mathcal{B}(t)$ Set of Byzantine workers $\delta_{1}$ $(\delta_{1},\delta_{2})-$heterogeneity parameter (2.4) $g_{i}^{t}$ Gradient from worker $i$ at iteration $t$$\cdot\|_{\infty}$ $\max\limits_{1\leqslant i\leqslant d}|\cdot|_{i}$
$B=B(t)$ Number of Byzantine workers $\delta_{2}$ $(\delta_{1},\delta_{2})-$heterogeneity parameter (2.4) $x^{*}$ Optimal solution of the objective function $\Delta_{d}^{1}$ $d$-dimensional simplex constraint on weights
$t$ Current iteration number $\sigma^{2}$ Variance of stochastic gradients (2.3) $\overline{x}^{t}$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}$ $\omega_{i}^{t}$ Weight assigned to worker $i$ at iteration $t$
$T$ Total number of iterations in training $l$ Local round length $\mathcal{KL}(p\|q)$ Kullback-Leibler divergence between distributions $p$ and $q$ $[\cdot]_{0}$ Non-negative projection: $\max\{0,\cdot\}$
[…]
Table 11: Notation Reference.
Appendix C General Inequalities and Lemmas
First, mention important inequalities that are used in further proofs. Consider a function $f$ satisfying Assumption 2.1, $g$ satisfying Assumptions 2.2(a) and $\varphi$ complying with Assumption 2.2(b). Then for any $i$ in the real numbers and for all vectors $x,y,x_{i}$ in $\mathbb{R}^{n}$ with a positive scalar $p$, the following inequalities hold.
$\left|\left\langle x,y\right\rangle\right|\quad$ $\leqslant\quad\frac{\|x\|^{2}}{2p}+\frac{p\|y\|^{2}}{2}$ (Young)
$\begin{split}-\left\langle x,y\right\rangle\quad&=\quad-\frac{\|x\|^{2}}{2}-\frac{\|y\|^{2}}{2}+\frac{\|x-y\|^{2}}{2}\\ \|x+y\|^{2}\quad&=\quad\|x\|^{2}+\|y\|^{2}+2\left\langle x,y\right\rangle\end{split}$ (Norm)
$\quad g(y)\quad$ $\geqslant\quad g(x)-\left\langle\nabla g(y),x-y\right\rangle-\frac{1}{2\mu}\|\nabla g(x)-\nabla g(y)\|^{2}$ ($\mu$-Conv)
$\begin{split}\quad\varphi(y)\quad&\geqslant\quad\varphi(x)+\left\langle\nabla\varphi(x),y-x\right\rangle\\ 0\quad&\geqslant\quad\left\langle\varphi(x)-\varphi(y),y-x\right\rangle\end{split}$ (Conv)
$\begin{split}\left\|\nabla f(x)-\nabla f(y)\right\|^{2}\quad&\leqslant\quad L^{2}\|x-y\|^{2}\\ \quad f(x)\quad&\leqslant\quad f(y)+\left\langle\nabla f(y),x-y\right\rangle+\frac{L}{2}\|x-y\|^{2}\\ f(x)\quad&\leqslant\quad f(y)-\left\langle\nabla f(x),y-x\right\rangle-\frac{1}{2L}\|\nabla f(x)-\nabla f(y)\|^{2}\end{split}$ (Lip)
$\begin{split}\left\|\sum_{i=1}^{n}x_{i}\right\|^{2}\quad&\leqslant\quad n\sum_{i=1}^{n}\|x_{i}\|^{2}\\ \left\|x+y\right\|^{2}\quad&\leqslant\quad(1+p)\|x\|^{2}+\left(1+\frac{1}{p}\right)\|y\|^{2}\end{split}$ (CS)
$\varphi\left(\frac{\sum_{i=1}^{n}w_{i}x_{i}}{\sum_{i=1}^{n}w_{i}}\right)\quad$ $\leqslant\quad\frac{\sum_{i=1}^{n}w_{i}\varphi(x_{i})}{\sum_{i=1}^{n}w_{i}}$ (Jen)
This section delineates a series of lemmas that form the cornerstone of our subsequent proofs. These lemmas encapsulate critical properties and bounds that are instrumental in establishing the theorems elaborated later in this paper.
The following lemma addresses a critical issue concerning the evaluation of the trial function and its deviation from $\nabla f_{1}$. As highlighted in the main part of our work, existing literature frequently overlooks the additional term in convergence that arises when employing a trial function. In this study, we rectify this oversight, thereby providing a more comprehensive understanding of the convergence behavior associated with trial functions.
Lemma C.1.
Suppose Assumption 2.1 holds. Then for all $x\in\mathcal{X}\subset\mathbb{R}^{d}$ with probability of at least $1-\widetilde{\delta}$ over a sample of size $N$, the following estimate, linking the trial function with the objective function on the server, is valid:
$\begin{split}\|\nabla f_{1}(x)-\nabla\hat{f}(x)\|_{2}^{2}~{}\leqslant~{}\zeta(N)=\widetilde{\mathcal{O}}\left(\frac{1}{N}\right),\end{split}$
Proof.
Given the norm inequality $\|\cdot\|_{2}\leqslant\sqrt{d}\cdot\|\cdot\|_{\infty}$, we can recast the scalar product in the following manner:
$\|\nabla f_{1}(x)-\nabla\hat{f}(x)\|_{2}^{2}$$\leqslant$ $d\cdot\|\nabla f_{1}(x^{t})-\nabla\hat{f}(x)\|_{\infty}^{2}.$ (4)
To establish the uniform convergence of $\left\|\nabla f_{1}(x)-\nabla\hat{f}(x)\right\|_{\infty}^{2}$, we employ Theorem 5 from (Shalev-Shwartz et al., 2009). This theorem provides a bound on the $\ell_{\infty}$-covering number of the function class =$\mathcal{F}=\{\xi\mapsto\nabla f_{1}(x;\xi)\mid x\in\mathcal{X}\}$. Given that $\mathcal{X}$ resides within an $\ell_{2}$-sphere, let us define it bound by $S$, the covering number for $\mathcal{X}$ using the Euclidean metric $d_{2}(x_{i},x_{j})=\|x_{i}-x_{j}\|_{2}$ is constrained as follows for $d>3$:
$N(\varepsilon,\mathcal{X},d_{2})=\mathcal{O}\left(d^{2}\left(\frac{S}{\varepsilon}\right)^{d}\right).$
In evaluating the covering numbers for $\mathcal{F}$ under the $\ell_{\infty}$ metric, where $\left\|\nabla f_{1}(x_{i};\cdot)-\nabla f_{1}(x_{j};\cdot)\right\|_{\infty}=\sup_{\xi}|\nabla f_{1}(x_{i};\xi)-\nabla f_{1}(x_{j};\xi)|$, the $L$-smoothness property facilitates the following assertion:
$\forall x_{i},x_{j}\in\mathcal{X}\hookrightarrow\left\|\nabla f_{1}(x_{i};\cdot)-\nabla f_{1}(x_{j};\cdot)\right\|_{\infty}\leqslant\left\|\nabla f_{1}(x_{i};\cdot)-\nabla f_{1}(x_{j};\cdot)\right\|_{2}\leqslant L\|x_{i}-x_{j}\|.$
This indicates that an $\varepsilon$-net for $\mathcal{X}$ in $d_{2}$ space concurrently serves as an $L\varepsilon$-net for $\mathcal{F}$ in $d_{\infty}$ space:
$N(\varepsilon,\mathcal{F},d_{\infty})\leqslant N(\varepsilon/L,\mathcal{X},d_{2})=\mathcal{O}\left(d^{2}\left(\frac{LS}{\varepsilon}\right)^{d}\right).$
Following this analysis, we derive an estimation consistent with the findings in (Shalev-Shwartz et al., 2009):
$\|\nabla f_{1}(x)-\nabla\hat{f}(x)\|_{\infty}^{2}=\widetilde{\mathcal{O}}\left(\frac{1}{N}\right).$
Defining the notation
[…]
and substituting this into (4) concludes the proof of the lemma.
This lemma is technical in nature, and we significantly benefit from the assertion established in the previous lemma. Ultimately, we derive an important estimate for the scalar product, which appears in many subsequent proofs throughout this work.
Lemma C.2.
Suppose Assumption 2.1 holds. Then for all $x\in\mathbb{R}^{d}$ and $g_{i}=g_{i}(x,\xi_{i})$, the following estimate is valid:
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$$\leqslant$ $-\frac{\gamma}{2}\left\|\nabla f(x)\right\|^{2}+\gamma\cdot\zeta(N)+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right).$
Proof.
We commence by examining the difference $\nabla f(x)-\nabla\hat{f}(x)$:
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle.$
Next, we continue with further manipulations on the first term:
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ (${\leqslant}$ $\frac{\gamma}{2}\left\|\nabla f(x)-\nabla\hat{f}(x)\right\|^{2}+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{1}(x)\right\|^{2}$
(${\leqslant}$ $\frac{\gamma}{2}\left(\left\|\nabla f(x)-\nabla f_{1}(x)\right\|^{2}+\left\|\nabla f_{1}(x)-\nabla\hat{f}(x)\right\|^{2}\right)$
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
[…]
and with the second term,
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ ${=}$ $-\frac{\gamma}{2}\|\nabla f(x)\|^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}(\nabla f_{i}(x)-\nabla f(x))\right\|^{2}$
(${\leqslant}$ $-\frac{\gamma}{2}\|\nabla f(x)\|^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
$+\frac{\gamma}{2G}\sum\limits_{i\in\mathcal{G}}\left\|(\nabla f_{i}(x)-\nabla f(x))\right\|^{2}$
[…]
$+\frac{\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x)\|^{2}\right).$
In summary, this supports the claim of the lemma. ∎
With this, we conclude the discussion of general statements. They are frequently used in our proofs in the upcoming sections of the Appendix. Next, we begin the examination of each method obtained individually.
Appendix D Proofs of BANT
In this section, we explore the theoretical underpinnings of the first proposed method, BANT. As outlined in Algorithm 1, this step diverges from the standard SGD approach primarily due to the distinct weight distribution that we subsequently allocate to the devices. Consequently, it is essential to conduct an analysis that takes this particular characteristic into account. To achieve final convergence rate, we demonstrate a supporting lemma that reinforces our findings.
Lemma D.1.
Under Assumptions 2.1, 2.2(b), 2.5, the following holds for the iteration of Algorithm 1:
$\hat{f}(x^{t+1})$$\leqslant$ $\hat{f}(x^{t})-\frac{\gamma\beta}{n}\left\langle\nabla\hat{f}(x^{t}),\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\|g_{i}^{t}\|^{2}.$
Proof.
Actually, the update step of the algorithm 1 is given by:
$x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\mathbb{I}_{[\theta_{i}^{t}>0]}\omega_{i}^{t}g_{i}^{t},$
where $g_{i}^{t}=g_{i}(x^{t},\xi_{i}^{t})$ and $\sum_{i=1}^{n}\omega_{i}^{t}=1$. Applying Jensen’s inequality for the convex function $\hat{f}$ (Assumption 2.2(b)) and denoting $\overline{\omega}_{i}^{t}=\frac{[\theta_{i}^{t}]_{0}}{\sum_{j=1}^{n}[\theta_{j}^{t}]_{0}}$:
$\hat{f}(x^{t+1})$ $=$ $\hat{f}\Bigl{(}\sum\limits_{i=1}^{n}\omega_{i}^{t}\left[x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right]\Bigr{)}$
$\leqslant$ $\sum\limits_{i=1}^{n}\omega_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$=$ $\sum\limits_{i\in\mathcal{B}}\omega_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)+\sum\limits_{i\in\mathcal{G}}\omega_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$\leqslant$ $\sum\limits_{i\in\mathcal{B}}(1-\beta)\omega_{i}^{t-1}\hat{f}(x^{t})+\sum\limits_{i\in\mathcal{B}}\beta\overline{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$+\sum\limits_{i\in\mathcal{G}}(1-\beta)\omega_{i}^{t-1}\hat{f}\left(x^{t}\right)+\sum\limits_{i\in\mathcal{G}}\beta\overline{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$=$ $(1-\beta)\hat{f}(x^{t})+\sum\limits_{i\in\mathcal{B}}\beta\overline{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)+\sum\limits_{i\in\mathcal{G}}\beta\overline{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right).$
In the inequality above, we make an estimation $\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)\leqslant\hat{f}\left(x^{t}\right)$, since the indicator guarantees us that we do not increase the trial function $\hat{f}$ by performing a step. By eliminating the weights $\omega_{i}^{t-1}$ accumulated from past iterations, we can rearrange the coefficients between Byzantine and honest workers in such a way that honest workers have higher weights. To achieve this, we sort the honest workers by increasing values of $\hat{f}$ and assign them coefficients $\omega_{i}$ in decreasing order. This permutation ensures that honest workers have higher weights and Byzantine workers have lower weights. This operation is valid because if $\overline{\omega}$ for some Byzantine worker is higher than for a honest worker, then this Byzantine has a greater influence on $\hat{f}$, and changing the weights would worsen the overall influence of these two workers. Therefore, with new weights $\{\widetilde{\omega}_{i}^{t}\}_{i=1}^{n}$:
$\hat{f}(x^{t+1})$$\leqslant$ $(1-\beta)\hat{f}(x^{t})+\sum\limits_{i\in\mathcal{B}}\beta\widetilde{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$+\sum\limits_{i\in\mathcal{G}}\beta\widetilde{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$\leqslant$ $(1-\beta)\hat{f}(x^{t})+\sum\limits_{i\in\mathcal{B}}\beta\widetilde{\omega}_{i}^{t}\hat{f}(x^{t})+\sum\limits_{i\in\mathcal{G}}\beta\widetilde{\omega}_{i}^{t}\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)$
$=$ $\hat{f}(x^{t})+(1-\beta)\left[\hat{f}(x^{t})-\hat{f}(x^{t})\right]+\sum\limits_{i\in\mathcal{B}}\beta\widetilde{\omega}_{i}^{t}\left[\hat{f}(x^{t})-\hat{f}(x^{t})\right]$
$+\sum\limits_{i\in\mathcal{G}}\beta\widetilde{\omega}_{i}^{t}\left[\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)-\hat{f}(x^{t})\right].$
Let us assign the coefficient $\nicefrac{{1}}{{n}}$ to all honest workers. This procedure is also valid. We sorted the weights and honest workers now have the greatest weights, thus, the sum of the coefficients of honest workers is at least $\nicefrac{{G}}{{n}}$. Moreover, the honest workers with the stronger influence have the greater weights which allows to equalize the total weight $\nicefrac{{G}}{{n}}$ between all $G$ workers. Thus, we get
$\textstyle{\hat{f}(x^{t+1})\leqslant\hat{f}(x^{t})+\frac{\beta}{n}\sum\limits_{i\in\mathcal{G}}\left[\hat{f}\left(x^{t}-\gamma\mathbb{I}_{[\theta_{i}^{t}>0]}g_{i}^{t}\right)-\hat{f}(x^{t})\right].}$
Now we can remove the indicator function because if $g_{i}^{t}$ minimizes the trial function, the indicator equals 1. If $g_{i}^{t}$ maximizes the trial function, the indicator excludes this gradient. However, we still account for it and maximize the trial function, thus:
$\hat{f}(x^{t+1})$$\leqslant$ $\hat{f}(x^{t})+\frac{\beta}{n}\sum\limits_{i\in\mathcal{G}}\left[\hat{f}\left(x^{t}-\gamma g_{i}^{t}\right)-\hat{f}(x^{t})\right]$
(${\leqslant}$ $\hat{f}(x^{t})+\frac{\beta}{n}\sum\limits_{i\in\mathcal{G}}\Biggl{[}\hat{f}(x^{t})-\gamma\left\langle\nabla\hat{f}(x^{t}),g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\|g_{i}^{t}\|^{2}-\hat{f}(x^{t})\Biggr{]}$
$=$ $\hat{f}(x^{t})-\frac{\gamma\beta}{n}\left\langle\nabla\hat{f}(x^{t}),\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\|g_{i}^{t}\|^{2}.$
We are now prepared to present the final result for the convex case. This theorem was introduced in the main part of our work, however, we will reiterate its formulation once more.
Theorem 3.1. Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n}{\beta G}+3\delta_{1}+6L\gamma\sigma^{2}+4\zeta(N).$
Proof.
According to the lemma (D.1):
$\hat{f}(x^{t+1})$$\leqslant$ $\hat{f}(x^{t})-\gamma\beta\left\langle\nabla\hat{f}(x^{t}),\frac{1}{n}\sum\limits_{i\in\mathcal{G}}\nabla g_{i}^{t}\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\left\|g_{i}^{t}\right\|^{2}.$
Taking the expectation of both sides of the inequality:
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\beta\cdot\frac{G}{n}\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
[…]
$+\frac{3\gamma\beta G}{2n}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma\beta G}{n}\zeta(N)-\frac{\gamma\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma\beta G}{2n}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta}{2n}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}\beta G}{2n}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\sigma^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}$
[…]
$+\frac{\gamma\beta G}{2n}(3+3L\gamma)\delta_{1}+\frac{3\gamma^{2}\beta LG}{2n}\sigma^{2}+\frac{\gamma\beta G}{n}\zeta(N).$
We first fix $\delta_{2}\leqslant\frac{1}{12}$. Then by choosing $\gamma\leqslant\frac{1}{13L}\leqslant\frac{1}{12L(1+\delta_{2})}$ and summing over the iterations, we get the bound:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n}{\beta G}+3\delta_{1}+6L\gamma\sigma^{2}+4\zeta(N).$
We have obtained the final statement of the theorem. From this, we can derive the convergence rate:
Corollary 3.2 Under the assumptions of Theorem 3.1, for solving the problem (1), after $T$ iterations with $\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]n}}{\sigma\sqrt{3LG\beta T}}\right\}$, the following holds:
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $=$ $\mathcal{O}\Biggl{(}\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}\quad+\frac{\sigma\cdot\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\Biggr{)};$
Proof of Corollary 3.2.
We proceed estimation, analogical to Lemma 4 from (Stich, 2019). Using the result of Theorem 3.1, we choose the appropriate $\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]n}}{\sigma\sqrt{3LG\beta T}}\right\}$. In that way, we obtain
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]52Ln}{\beta GT}+\frac{2\sigma\sqrt{6\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}}{\sqrt{\beta GT}}+3\delta_{1}+2\zeta(N),$
that ends the proof. ∎
We have obtained the result for the convex case. However, we wish to extend the theory to other cases. Using the obtained result, let us proceed to the $\mu$-strongly convex case and derive an estimate for it.
Theorem D.2.
Under Assumptions 2.1, 2.2(a), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]$$\leqslant$ $\left(1-\frac{\gamma\beta G\mu}{4n}\right)^{t}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+\frac{3}{\mu}\delta_{1}+\frac{6L\gamma}{\mu}\sigma^{2}+\frac{2}{\mu}\zeta(N).$
Proof.
From Theorem 3.1, we have
$\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4n}{\gamma\beta G}\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{t+1})\right]+3\delta_{1}+6L\gamma\sigma^{2}+4\zeta(N).$
Let us examine the left-hand side of the inequality in more detail:
$\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}$ $=$ $\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}+\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}-\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}$
(${\geqslant}$ $\frac{1}{2}\mathbb{E}\left\|\nabla f(x^{t})+\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}-\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}$
[…]
Returning to the initial inequality,
$\left(\frac{1}{2}-2\delta_{2}\right)\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla\hat{f}(\hat{x}^{*})\right\|^{2}$$\leqslant$ $\frac{4n}{\gamma\beta G}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+5\delta_{1}+6L\gamma\sigma^{2}+6\zeta(N)$
Taking into account that $\delta_{2}\leqslant\frac{1}{12}$ and due to ($\mu$-Conv), we get:
$\underbrace{\frac{1}{2}\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla\hat{f}(\hat{x}^{*})\right\|^{2}}_{\geqslant~{}\mu\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]}$$\leqslant$ $\frac{6n}{\gamma\beta G}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+8\delta_{1}+9L\gamma\sigma^{2}+9\zeta(N)$
$\frac{6n}{\gamma\beta G}\mathbb{E}\hat{f}(x^{t+1})-\mu\mathbb{E}\hat{f}(\hat{x}^{*})$$\leqslant$ $\left(\frac{6n}{\gamma\beta G}-\mu\right)\mathbb{E}\hat{f}(x^{t})+8\delta_{1}+9L\gamma\sigma^{2}+9\zeta(N).$
Defining $\gamma^{\prime}=\frac{\gamma\beta G}{6n},$
$\mathbb{E}\hat{f}(x^{t+1})-\gamma^{\prime}\mu\mathbb{E}\hat{f}(\hat{x}^{*})$$\leqslant$ $\left(1-\gamma^{\prime}\mu\right)\mathbb{E}\hat{f}(x^{t})+8\gamma^{\prime}\delta_{1}+9\gamma^{\prime}L\gamma\sigma^{2}+9\gamma^{\prime}\zeta(N).$
We add to both sides of inequality the term $(-1+\gamma^{\prime}\mu)\mathbb{E}\hat{f}(\hat{x}^{*})$:
$\left[\mathbb{E}\hat{f}(x^{t+1})-\mathbb{E}\hat{f}(\hat{x}^{*})\right]$$\leqslant$ $\left(1-\gamma^{\prime}\mu\right)\left[\mathbb{E}\hat{f}(x^{t})-\mathbb{E}\hat{f}(\hat{x}^{*})\right]+8\gamma^{\prime}\delta_{1}+\frac{9L\gamma\gamma^{\prime}}{G}\sigma^{2}+9\gamma^{\prime}\zeta(N).$
Applying this inequality to the first term on the right side $t$ times, we obtain:
$\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]$$\leqslant$ $\left(1-\gamma^{\prime}\mu\right)^{t}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+\underbrace{\sum\limits_{i=0}^{t-1}\left(1-\gamma^{\prime}\mu\right)^{i}}_{\leqslant\frac{1}{\gamma^{\prime}\mu}}\left[8\gamma^{\prime}\delta_{1}+9L\gamma\gamma^{\prime}\sigma^{2}+9\gamma^{\prime}\zeta(N)\right].$
$\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]$$\leqslant$ $\left(1-\gamma^{\prime}\mu\right)^{t}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+\frac{8}{\mu}\delta_{1}+\frac{9L\gamma}{\mu}\sigma^{2}+\frac{9}{\mu}\zeta(N).$
Similarly, from this estimate, we derive the final convergence rate for the $\mu$-strongly convex setting.
Corollary D.3.
Under the assumptions of Theorem D.2, for solving the problem (1), after $T$ iterations with special tunings of $\gamma$:
$\mathbb{E}\left[\hat{f}(x^{T})-\hat{f}(\hat{x}^{*})\right]$ $=$ $\mathcal{\widetilde{O}}\bigg{(}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\exp\left[-\frac{\mu\beta GT}{4Ln}\right]+\frac{Ln}{\mu^{2}\beta GT}\sigma^{2}+\frac{1}{\mu}\delta_{1}+\frac{1}{\mu}\zeta(N)\bigg{)}.$
Proof.
In Theorem D.2, we obtain classic result for SGD. We use Lemma 2 from (Stich, 2019) and appropriate special tunings of $\gamma$:
$\gamma$$\leqslant$ $\min\left\{\frac{1}{13L},\frac{4n\log\left(\max\left\{2,\frac{\mu^{2}\beta G\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{36Ln\sigma_{*}^{2}}\right\}\right)}{\mu\beta GT}\right\}.$
We obtain the final convergence. ∎
With this, we conclude this section. In summary, we examine the proof for both the convex and strongly convex cases and obtain the final convergence estimates.
Appendix E Proofs of AutoBANT
Now, let us turn our attention to the second of our methods. As previously mentioned, BANT has certain moments to be discussed. It is important to note the adverse impact of the parameter $\beta$. The introduction of this momentum term aimed to protect honest clients from rapidly decreasing their trust scores due to unfavorable stochastic gradients, but it inadvertently enables Byzantine agents to maintain their weights despite their attacks. To address this issue, we add an indicator to the algorithm for detecting Byzantine devices, although this limits its theoretical part in common non-convex scenarios.
Our goal is to learn how to circumvent this limitation. To achieve this, we tackle an additional subproblem related to weight assignment. This step represents a key distinction in our theoretical analysis. We can assert that we are solving this weight distribution minimization subproblem with a certain error margin $\delta$, which will be reflected in the final convergence results. Next, we present the theorem discussed in the main part of the paper, along with its complete proof.
Theorem 3.3. Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
Proof.
The iterative update formula for $x^{t+1}$ is given by
$x^{t+1}$ $=$ $x^{t}-\gamma\sum_{i=1}^{n}\left(\arg\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i=1}^{n}\omega_{i}g_{i}^{t}\right]\right)g_{i}^{t},$
which leads to an upper bound on $\hat{f}(x^{t+1})$:
$\hat{f}(x^{t+1})$$\leqslant$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i=1}^{n}\omega_{i}g_{i}^{t}\right]+\delta$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
Taking the expectation,
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
[…]
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f(x^{t})\right\|^{2}+\delta.$
Due to the fact that $\mathbb{E}g_{i}^{t}=\nabla f_{i}(x^{t})$ and $\mathbb{E}\langle\nabla f_{i}(x^{t})-g_{i}^{t},\nabla f_{j}(x^{t})-g_{j}^{t}\rangle=0$,
$\mathbb{E}\hat{f}(x^{t+1})$ (${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left(\frac{1}{G}\sum_{i\in\mathcal{G}}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\frac{1}{G^{2}}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
[…]
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
$=$ $\mathbb{E}[\hat{f}(x^{t})]-\frac{\gamma}{2}\left[1-3L\gamma-(3+3L\gamma)\delta_{2}\right]\|\nabla f(x^{t})\|^{2}$
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)+\delta.$
We first fix $\delta_{2}\leqslant\frac{1}{12}$. By choosing $\gamma\leqslant\frac{1}{13L}\leqslant\frac{1}{12L(1+\delta_{2})}$, and summing over the iterations, we get
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
We have successfully proven the obtained result. Let us also recall that we formulated the final estimate in Corollary 3.4. We will omit the proof of Corollary 3.4 since it entirely replicates the proof of Corollary 3.2.
With this, we conclude our proof of the foundational versions of the algorithms. We establish all the formulated statements and derive convergence estimates for the strongly convex, convex, and non-convex cases. The subsequent sections of the Appendix are dedicated to exploring extensions that hold significant importance in our study.
Appendix F Scaled methods
In this section, we provide a detailed analysis of our Byzantine-robust methods extended to adaptive methods, as mentioned in the main part. Specifically, we consider the application of our techniques to methods like Adam and RMSProp. We describe the formal description of Scaled BANT and Scaled AutoBANT methods, which utilize a diagonal preconditioner $(\hat{P}^{t})^{-1}$, which scales a gradient to $(\hat{P}^{t})^{-1}g_{i}^{t}$, and the step is performed using this scaled gradient. From iteration to iteration, the matrix $P^{t}$ changes, e.g. the following rule can be used.
$(P^{t})^{2}=\beta_{t}(P^{t-1})^{2}+(1-\beta_{t})(H^{t})^{2}.$ (5)
This update scheme is satisfied by Adam-based methods with $(H^{t})^{2}=\text{diag}(g^{t}\odot g^{t})$ and by AdaHessian (Yao et al., 2021) with $(H^{t})^{2}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))^{2}$, where $\odot$ denotes the component-wise product between two vectors, and $z^{t}$ are from Rademacher distribution, i.e. all components from vector are independent and equal to ±$\pm 1$ with probability $\nicefrac{{1}}{{2}}$.
We want the preconditioner being a positive define matrix, thus, it is typical to modify $P^{t}$ a bit:
$(\hat{P}^{t})_{ii}=\max\{e,|P^{t}|_{ii}\},$ (6)
where $e$ is a (small) positive parameter. There are also other possible update rules, one of which is
$P^{t}=\beta_{t}(P^{t-1})+(1-\beta_{t})H^{t}.$
For example, such a rule is extended in OASIS (Jahani et al., 2022) with $\beta_{t}\equiv\beta$ and $H^{t}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))$. We also note additional details in the construction of a positively defined preconditioner. We can also alternatively define $\hat{P}^{t}$ as $(\hat{P})_{ii}=|P|_{ii}+e$. Most importantly, both of these approaches construct diagonal matrices with positive elements. We introduce the crucial for our analysis assumption.
Assumption F.1.
For any $t\geqslant 1$, we have $\alpha I\preccurlyeq\hat{P}^{t}\preccurlyeq\Gamma I$.
The correct proof of this statement, as well as a more detailed description of the diagonal preconditioner, is provided in (Sadiev et al., 2024). We mention that for AdaHessian and OASIS preconditioners, $\Gamma=\sqrt{d}L$, and for Adam and RMSProp, under the condition $\|\nabla f(x)\|\leqslant M,\Gamma=M$. Thus, having constructed a diagonal preconditioner, we can proceed to the analysis of scaled methods.
F.1  Scaled BANT
Before presenting the results for Scaled BANT (Algorithm 5), we need to provide the preliminary analysis. To derive the final estimation, let us prove auxiliary lemmas.
Lemma F.2.
If the diagonal preconditioner $\hat{P}$ is such that $\alpha I\preccurlyeq\hat{P}\preccurlyeq\Gamma I$, the following estimates are valid:
  1. (a)
    $\frac{1}{\Gamma}\|g\|^{2}\leqslant$ $\left\|g\right\|_{\hat{P}^{-1}}^{2}$ $\leqslant\frac{1}{\alpha}\|g\|^{2},$
  2. (b)
    $\frac{1}{\Gamma^{2}}\|g\|^{2}\leqslant$ $\left\|\hat{P}^{-1}g\right\|^{2}$ $\leqslant\frac{1}{\alpha^{2}}\|g\|^{2},$
where $\langle h,g\rangle_{\hat{P}^{-1}}\overset{def}{=}\left\langle h,\hat{P}^{-1}g\right\rangle$, $h,g\in\mathbb{R}^{d}$
Proof.
$I\preccurlyeq\frac{1}{\alpha}\hat{P}~{}\Rightarrow~{}\left\|\hat{P}^{-1}g\right\|^{2}=\left\langle I\hat{P}^{-1}g,\hat{P}^{-1}g\right\rangle$$\leqslant$ $\frac{1}{\alpha}\left\langle g,\hat{P}^{-1}g\right\rangle~{}\overset{def}{=}~{}\frac{1}{\alpha}\left\langle g,g\right\rangle_{\hat{P}^{-1}}=\frac{1}{\alpha}\|g\|_{\hat{P}^{-1}}^{2}.$
$\left\langle g,\hat{P}^{-1}g\right\rangle~{}$$\leqslant$ $~{}\frac{1}{\alpha}\left\langle g,g\right\rangle=\frac{1}{\alpha}\|g\|^{2}.$
$\hat{P}\preccurlyeq\Gamma I~{}\Rightarrow~{}\left\|\hat{P}^{-1}g\right\|^{2}=\left\langle I\hat{P}^{-1}g,\hat{P}^{-1}g\right\rangle$$\geqslant$ $\frac{1}{\Gamma}\left\langle g,\hat{P}^{-1}g\right\rangle~{}\overset{def}{=}~{}\frac{1}{\Gamma}\left\langle g,g\right\rangle_{\hat{P}^{-1}}=\frac{1}{\Gamma}\|g\|_{\hat{P}^{-1}}^{2}.$
$\left\langle g,\hat{P}^{-1}g\right\rangle~{}$$\geqslant$ $~{}\frac{1}{\Gamma}\left\langle g,g\right\rangle=\frac{1}{\Gamma}\|g\|^{2}.$
Algorithm 5 Scaled BANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$
2:  Parameters: Stepsize $\gamma>0$, momentum parameter $\beta\in[0,1]$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Server sends $x^{t}$ to each worker
5:     for all workers $i=0,1,2,\ldots,n$ in parallel do
6:        Generate $\xi_{i}^{t}$ independently
7:        Compute stochastic gradient $g_{i}(x^{t},\xi_{i})$
8:        Send $g_{i}^{t}=g_{i}(x^{t},\xi_{i})$ to server
9:     end for
10:     $\omega^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})]_{0}}{\sum\limits_{j=1}^{n}[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{j}^{t})]_{0}}$
11:     if each $\left[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})\right]_{0}=0$ then
12:        $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{1}{n}$
13:     end if
14:     $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\mathbb{I}_{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})>0]}\omega_{i}^{t}g_{i}^{t}$
15:     $\hat{P}^{t}$ is the function of $\hat{P}^{t-1}$ and $H^{t}$, e.g., as (5) + (6)
16:  end for
17:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
Lemma F.3 (Scaled version of Lemma D.1).
Under Assumptions 2.1, 2.2(b), 2.5, F.1, the following holds for the iteration of Algorithm 5:
$\hat{f}(x^{t+1})$$\leqslant$ $\hat{f}(x^{t})-\gamma\beta\frac{1}{n}\sum\limits_{i\in\mathcal{G}}\left\langle\nabla\hat{f}(x^{t}),(\hat{P}^{t})^{-1}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\|(\hat{P}^{t})^{-1}g_{i}^{t}\|^{2}$
Proof.
Our analysis implies the same as was done in Lemma D.1. Since the only thing that has changed is that we additionally scale the gradient at each step, the analysis is similar and we obtain the final estimate. ∎
Lemma F.4 (Scaled version of Lemma C.2).
Suppose Assumption 2.1 holds. Then for all $x\in\mathbb{R}^{d}$ and $g_{i}=g_{i}(x,\xi_{i})$, the following estimate is valid:
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle^{2}$$\leqslant$ $-\frac{\gamma}{2\Gamma}\left\|\nabla f(x)\right\|^{2}+\frac{\gamma}{2\alpha}\zeta(N)+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right).$
Proof.
We commence by examining the difference $\nabla f(x)-\nabla\hat{f}(x)$:
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}.$
Next, we continue with further manipulations on the first term:
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ (${\leqslant}$ $\frac{\gamma}{2}\left\|\nabla f(x)-\nabla\hat{f}(x)\right\|_{\hat{P}^{-1}}^{2}+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
[…]
and with the second term,
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ ${=}$ $-\frac{\gamma}{2}\|\nabla f(x)\|_{\hat{P}^{-1}}^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}(\nabla f_{i}(x)-\nabla f(x))\right\|_{\hat{P}^{-1}}^{2}$
(${\leqslant}$ $-\frac{\gamma}{2}\|\nabla f(x)\|_{\hat{P}^{-1}}^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
$+\frac{\gamma}{2G}\sum\limits_{i\in\mathcal{G}}\left\|(\nabla f_{i}(x)-\nabla f(x))\right\|_{\hat{P}^{-1}}^{2}$
[…]
$+\frac{\gamma}{2\alpha G}\sum\limits_{i\in\mathcal{G}}\left\|(\nabla f_{i}(x)-\nabla f(x))\right\|^{2}$
[…]
$+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x)\|^{2}\right).$
Summing up substantiates the claim of the lemma. ∎
We are now ready to write out the main results for the scaled methods.
Theorem F.5.
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{2\Gamma-\alpha}{\alpha+\frac{4\Gamma^{2}}{\alpha^{2}}}$, 2.5, F.1, for solving the problem (1), after $T$ iteration of Algorithm 5 with $\gamma\leqslant\frac{\alpha}{12L}$, it holds that
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}\leqslant\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{n\Gamma}{\beta G}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N).$
Proof of Theorem F.5.
Similar to the aforementioned Lemma D.1 (with the preconditioner added):
$\hat{f}(x^{t+1})$$\leqslant$ $\hat{f}(x^{t})-\gamma\beta\frac{1}{n}\sum\limits_{i\in\mathcal{G}}\left\langle\nabla\hat{f}(x^{t}),(\hat{P}^{t})^{-1}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}\beta}{2n}\sum\limits_{i\in\mathcal{G}}\|(\hat{P}^{t})^{-1}g_{i}^{t}\|^{2}$
[…]
Taking the expectation of both sides of the inequality:
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\beta\cdot\frac{G}{n}\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}+\frac{L\gamma^{2}\beta}{2n\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
[…]
$+\frac{\gamma\beta G}{2n\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}\beta}{2n\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma\beta G}{2n\alpha}\zeta(N)-\frac{\gamma\beta G}{2n\Gamma}\left\|\nabla f(x^{t})\right\|^{2}+\frac{\gamma\beta G}{2n\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta}{2n\alpha^{2}}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}\beta G}{2n\alpha^{2}}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\sigma^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
[…]
$+\frac{\gamma\beta G}{2n\alpha}(1+\frac{3L\gamma}{\alpha})\delta_{1}+\frac{3\gamma^{2}\beta LG}{2n\alpha^{2}}\sigma^{2}+\frac{\gamma\beta G}{2n\alpha}\zeta(N).$
Now we have to choose $\gamma$. We want $\left[1-3L\gamma\frac{\Gamma}{\alpha^{2}}-(1+3L\gamma)\frac{\Gamma}{\alpha^{2}}\delta_{2}\right]\geqslant\frac{1}{2}$. Then
$\gamma$$\leqslant$ $\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{6L\frac{\Gamma}{\alpha^{2}}(1+\delta_{2})}.$
Let us choose $\delta_{2}\leqslant\frac{2\Gamma-\alpha}{\alpha+\frac{4\Gamma^{2}}{\alpha^{2}}}$, then we have $\frac{1}{2}\leqslant\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{\frac{\Gamma}{\alpha}(1+\delta_{2})}$. Thus,
$\gamma$$\leqslant$ $\frac{\alpha}{12L}~{}~{}\leqslant~{}~{}\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{6L\frac{\Gamma}{\alpha^{2}}(1+\delta_{2})}.$
Using $\gamma\leqslant\frac{\alpha}{12L}$ and summing over the iterations, we get the bound:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n\Gamma}{\beta G}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N).$
Now we provide the final convergence rate.
Corollary F.6.
Under assumptions of Theorem F.5 for solving the problem (1), after $T$ iterations of Algorithm F.1 with special tunings of $\gamma$, the following holds:
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\|\nabla f(x^{t})\|^{2}$ $=$ $\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}\cdot\frac{\Gamma}{\alpha}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}}{\sqrt{\beta GT}}\cdot\frac{\Gamma}{\alpha}+\left(\delta_{1}+\zeta(N)\right)\cdot\frac{\Gamma}{\alpha}\right).$
The proof of Corollary F.6 completely mirrors the proof of Corollary 3.2.
Remark F.7.
We get a result similar to Corollary 3.2, but with the deterioration that each term is multiplied by an additional constant $\nicefrac{{\Gamma}}{{\alpha}}>1$. This result suits us, since in (Sadiev et al., 2024) the result for the Scaled SARAH method corresponds similarly to the result for the classical SARAH (Nguyen et al., 2017) method.
F.2  Scaled AutoBANT
Now, let us consider the second algorithm for scaled methods - Scaled AutoBANT (Algorithm 6). This section presents an algorithm which is an adaptive version of Algorithm 2 taking into account the diagonal preconditioner. Now we provide an estimate for the convergence of the Scaled AutoBANT method.
Algorithm 6 Scaled AutoBANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$
2:  Parameters: Stepsize $\gamma>0$, error accuracy $\delta$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Server sends $x^{t}$ to each worker
5:     for all workers $i=0,1,2,\ldots,n$ in parallel do
6:        Generate $\xi_{i}^{t}$ independently
7:        Compute stochastic gradient $g_{i}(x^{t},\xi_{i})$
8:        Send $g_{i}^{t}=g_{i}(x^{t},\xi_{i})$ to server
9:     end for
10:     $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:     $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
12:     $\hat{P}^{t}$ is the function of $\hat{P}^{t-1}$ and $H^{t}$, e.g., as (5) + (6)
13:  end for
14:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
Theorem F.8.
Under Assumptions 2.1, 2.2(c) 2.3, 2.4 with $\delta_{2}\leqslant 0.25$, 2.5, F.1, for solving the problem (1), after $T$ iterations of Algorithm 6 with $\gamma\leqslant\frac{\alpha}{12L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N)+4\Gamma\frac{\delta}{\gamma}.$
Proof of Theorem F.8.
Note we estimate the trial function value:
$\hat{f}(x^{t+1})$$\leqslant$ $\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)+\delta$
$\leqslant$ $\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right)+\delta$
(${\leqslant}$ $\hat{f}\left(x^{t}\right)-\gamma\left\langle\nabla\hat{f}(x^{t}),(\hat{P}^{t})^{-1}\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|(\hat{P}^{t})^{-1}\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
[…]
Taking the expectation of both sides of the inequality:
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
[…]
$+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma}{2\alpha}\zeta(N)-\frac{\gamma}{2\Gamma}\left\|\nabla f(x^{t})\right\|^{2}+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta.$
Due to the fact that $\mathbb{E}g_{i}^{t}=\nabla f_{i}(x^{t})$ and $\mathbb{E}\langle\nabla f_{i}(x^{t})-g_{i}^{t},\nabla f_{j}(x^{t})-g_{j}^{t}\rangle=0$,
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma}{2\alpha}\zeta(N)-\frac{\gamma}{2\Gamma}\left\|\nabla f(x^{t})\right\|^{2}+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2G\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2G^{2}\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
[…]
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{1}{G}\sigma^{2}\right)+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
[…]
$+\frac{\gamma}{2\alpha}(1+\frac{3L\gamma}{\alpha})\delta_{1}+\frac{3L\gamma^{2}}{2\alpha^{2}G}\sigma^{2}+\frac{\gamma}{2\alpha}\zeta(N)+\delta.$
Now we have to choose $\gamma$: We want $\left[1-3L\gamma\frac{\Gamma}{\alpha^{2}}-(1+3L\gamma)\frac{\Gamma}{\alpha^{2}}\delta_{2}\right]\geqslant\frac{1}{2}$. Then
$\gamma$$\leqslant$ $\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{6L\frac{\Gamma}{\alpha^{2}}(1+\delta_{2})}.$
Let us choose $\delta_{2}\leqslant\frac{2\Gamma-\alpha}{\alpha+\frac{4\Gamma^{2}}{\alpha^{2}}}$, then we have $\frac{1}{2}\leqslant\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{\frac{\Gamma}{\alpha}(1+\delta_{2})}$. Thus,
$\gamma$$\leqslant$ $\frac{\alpha}{12L}~{}~{}\leqslant~{}~{}\frac{1-\frac{2\Gamma}{\alpha^{2}}\delta_{2}}{6L\frac{\Gamma}{\alpha^{2}}(1+\delta_{2})}.$
Using $\gamma\leqslant\frac{\alpha}{12L}$, summing over the iterations and taking the expected value at the initial point and at the optimum point, we get the bound:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N)+4\Gamma\frac{\delta}{\gamma}.$
Now, let us present the final convergence rate for this algorithm.
Corollary F.9.
Under assumptions of Theorem F.8, for solving the problem (1), after $T$ iterations of Algorithm (6) with special tunings of $\gamma$, the following holds:
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\|\nabla f(x^{t})\|^{2}$ $=$ $\mathcal{O}\Biggl{(}\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}{T}\cdot\frac{\Gamma}{\alpha}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}}{\sqrt{T}}\cdot\frac{\Gamma}{\alpha}+\left(\delta_{1}+\zeta(N)\right)\cdot\frac{\Gamma}{\alpha}$
$+\delta\Biggl{(}L+\frac{\sqrt{TL}\sigma}{\sqrt{\mathbb{E}[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})]G}}\Biggr{)}\cdot\frac{\Gamma}{\alpha}\Biggr{)}$
Remark F.10.
The boundary is the same as for AutoBANT, with the only aggravation that several summands are multiplied by $\frac{\Gamma}{\alpha}>1$. This result suits us for the same reason as the result of the Scaled BANT method.
The proof of Corollary F.9 completely mirrors the proof of Corollary 3.4.
Appendix G Local methods
As highlighted in the main section, the significant expense associated with communication remains a critical concern in various fields. The communication bottleneck can act as a substantial barrier, limiting efficiency and hindering progress. To address this challenge, many researches are turning to local approaches, which focus on minimizing the need for exchanging the information (Woodworth et al., 2020; Koloskova et al., 2020; Khaled et al., 2020; Gorbunov et al., 2021b). The idea is that each device performs a predefined number of local steps without utilizing information from other devices, and at the end of such a round, the server performs a mutual update. In this section we adapt our AutoBANT algorithm to this scenario. Below we present the formal description of the Local AutoBANT method (Algorithm 7).
Algorithm 7 Local AutoBANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$, local round length $l$
2:  Parameters: Stepsize $\gamma>0$, error accuracy $\delta$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     if $t=0$ then
5:        Server sends $x^{0}$ to each worker
6:     end if
7:     for all workers $i=0,1,2,\ldots,n$ in parallel do
8:        Generate $\xi_{i}^{t}$ independently
9:        Compute stochastic gradient $g_{i}^{t}=g_{i}(x^{t},\xi_{i})$
10:        if $t\neq t_{k\cdot l}\left(\text{for some~{}}k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}\right)\wedge t\neq T-1$ then
11:           $x_{i}^{t+1}=x_{i}^{t}-\gamma g_{i}^{t}$
12:        else
13:           Send $x_{i}^{t}-\gamma g_{i}^{t}$ to server
14:        end if
15:     end for
16:     if $t=t_{k\cdot l}\left(\text{for some~{}}k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}\right)\vee t=T-1$ then
17:        $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(\sum\limits_{i=1}^{n}\omega_{i}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)\right)$
18:        $x^{t+1}=\sum\limits_{i=1}^{n}\omega_{i}^{t}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)$
19:        Server sends $x^{t+1}$ to each worker
20:     end if
21:  end for
22:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
In the convergence analysis of Algorithm 7, we assume that at each local round, at least one device (including the server) acts as an honest worker. This implies that this device computes an honest stochastic gradient at each iteration of the round. This requirement is a natural extension of the assumption made in the analysis of the basic version of our methods, where we required at least one honest device (including the server) at each iteration (or in a local round of length 1).
The analysis has some specific details. The key component is estimating how far devices can "move apart" from each other during a local round. We begin with this estimation and present the following lemma.
Lemma G.1.
Under Assumptions 2.1, 2.2(b), 2.3, 2.4, 2.5, at each iteration $t$ of Algorithm 7 with $\gamma\leqslant\frac{1}{4(l-1)L}$, the following estimate is valid:
$\mathbb{E}V^{t}\leqslant\frac{9\delta_{2}\gamma}{2L}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+\frac{9\delta_{1}\gamma}{2L}(l-1)+3\gamma^{2}\sigma^{2}(l-1),$
where $t_{k\cdot l}$ for some $k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$ is the past to $t$-th iteration aggregation round, $V^{t}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$ and $\overline{x}^{t}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}$.
Proof.
Utilizing notation $V^{t}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$ and $\overline{x}^{t}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}$ we mention, that for all iterations $t+1$, such that $t+1=t_{k\cdot l}$ we have $V^{t+1}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|x_{i}^{t+1}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t+1}\right\|^{2}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|x^{t+1}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x^{t+1}\right\|^{2}=0$. For the rest iterations we write the step of the local update and use (Norm):
$\mathbb{E}\left\|x_{i}^{t+1}-\overline{x}^{t+1}\right\|^{2}$ $=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\gamma\mathbb{E}\left\langle x_{i}^{t}-\overline{x}^{t},g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle$
$=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\gamma\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle.$
Taking average over $i\in\mathcal{G}$,
$\mathbb{E}V^{t+1}$ $=$ $\mathbb{E}V^{t}+\frac{\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+2\gamma\left\langle\overline{x}^{t}-\overline{x}^{t},\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$ (7)
$=$ $\mathbb{E}V^{t}+\frac{\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle.$
Now we need to estimate the second term. We start with (Norm):
$\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$ $=$ $\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle$
=$\overset{(i)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\nabla f_{i}(x_{i}^{t}),\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle$
=$\overset{(ii)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle,$
where (i) was made (Norm), applied to the first norm and (ii) by taking expectation of the first scalar product and obtaining it equal to zero. Next, averaging over $i\in\mathcal{G}$ and transforming the scalar product,
$\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$ $=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$ (8)
$+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$\leqslant$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$\overset{(i)}{\leqslant}$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\sigma^{2},$
where (i) was made according to Assumption 2.3. To estimate the norm, we again use (Norm):
$\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$ $=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\langle\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t}),\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
$=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$-\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f(\overline{x}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
[…]
(${\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
$\overset{(i)}{\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+2(\delta_{1}+\delta_{2}\left\|\nabla f(\overline{x}^{t})\right\|^{2})$
(${\leqslant}$ $\frac{2}{G}\sum\limits_{i\in\mathcal{G}}\left(2L\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})-\left\langle\nabla f_{i}(x_{i}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle\right)\right)$
$+2(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2}),$
where (i) was made according to Assumption 2.4. Combining it with (7) and (8),
$\mathbb{E}V^{t+1}$$\leqslant$ $\mathbb{E}V^{t}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)$
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
$=$ $\mathbb{E}V^{t}-\frac{2\gamma(1-2L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)$
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}.$
Taking $\gamma\leqslant\frac{1}{4L}$ and applying (Conv) to scalar product,
$\mathbb{E}V^{t+1}$$\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+\frac{4L\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$ (9)
$=$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(f_{i}(\overline{x}^{t})-f_{i}(x_{i}^{t})\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
(${\leqslant}$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\left\langle\nabla f_{i}(x_{i}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
[…]
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
$\leqslant$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\left\langle\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\left\langle\nabla f(\overline{x}^{t}),\overline{x}^{t}-x_{i}^{t}\right\rangle+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}\right)$
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2}$
$\overset{(i)}{\leqslant}$ $\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{G}\sum\limits_{i\in\mathcal{G}}\left(\frac{1}{2L}\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\|^{2}+L\|\overline{x}^{t}-x_{i}^{t}\|^{2}+\frac{L}{2}\|\overline{x}^{t}-x_{i}^{t}\|^{2}+f(\overline{x}^{t})-f(x_{i}^{t})\right)$
$+2\gamma^{2}(\delta_{1}+\delta_{2}\|\nabla f(\overline{x}^{t})\|^{2})+\gamma^{2}\sigma^{2},$
where (i) was made with applying (Young) to the first scalar product and applying (Lip) to second one. Using (Jen) we derive $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left(f(\overline{x}^{t})-f(x_{i}^{t})\right)=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left(f\left(\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}\right)-f(x_{i}^{t})\right)\leqslant\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left(\frac{1}{G}\sum\limits_{i\in\mathcal{G}}f(x_{i}^{t})-f(x_{i}^{t})\right)=0$. In that way, we proceed (9) using this fact and taking expectation together with applying Assumption 2.4:
$\mathbb{E}V^{t+1}$$\leqslant$ $\mathbb{E}V^{t}+\frac{3L\gamma(1+4L\gamma)}{2}\mathbb{E}V^{t}+\frac{\gamma(1+4L\gamma)}{2L}\delta_{2}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\frac{\gamma(1+4L\gamma)}{2L}\delta_{1}$
$+2\gamma^{2}\left(\delta_{1}+\delta_{2}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}\right)+\gamma^{2}\sigma^{2}$
Going into recursion up to the past aggregation round, which was on the iteration $t_{k\cdot l}$ for some $k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$, together with using $\gamma\leqslant\frac{1}{4L}$ choice, we obtain
$\mathbb{E}V^{t}$$\leqslant$ $(1+3L\gamma)\mathbb{E}V^{t-1}+\delta_{2}\gamma\left(2\gamma+\frac{1}{L}\right)\mathbb{E}\left\|\nabla f(\overline{x}^{t-1})\right\|^{2}+\delta_{1}\gamma\left(2\gamma+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}$
$\leqslant$ $(1+3L\gamma)^{t-t_{k\cdot l}}\mathbb{E}V^{t_{k\cdot l}}+\delta_{2}\gamma\left(2\gamma+\frac{1}{L}\right)\sum\limits_{j=t_{k\cdot l}}^{t-1}(1+3L\gamma)^{j-t_{k\cdot l}}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}$
$+\left(\delta_{1}\gamma\left(2\gamma+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)\sum\limits_{j=t_{k\cdot l}}^{t-1}\left(1+3L\gamma\right)^{j-t_{k\cdot l}}$
$\leqslant$ $\delta_{2}\gamma\left(2\gamma+\frac{1}{L}\right)(1+3L\gamma)^{l-1}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+\left(\delta_{1}\gamma\left(2\gamma+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)(1+3L\gamma)^{l-1}.$
Now we tune $\gamma\leqslant\frac{1}{4(l-1)L}$. Note it is smallest of all previous $\gamma$, since $l\geqslant 2$ and consequently all previous transitions hold true. In that, way, using $\left(1+3L\gamma\right)^{l-1}\leqslant\left(1+\frac{3}{4(l-1)}\right)^{l-1}\leqslant\left(1+\frac{1}{l-1}\right)^{l-1}\leqslant 3$,
$\mathbb{E}V^{t}$$\leqslant$ $3\delta_{2}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+3\left(\delta_{1}\gamma\left(\frac{1}{2(l-1)L}+\frac{1}{L}\right)+\gamma^{2}\sigma^{2}\right)(l-1)$
$\leqslant$ $\frac{9\delta_{2}\gamma}{2L}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+\frac{9\delta_{1}\gamma}{2L}(l-1)+3\gamma^{2}\sigma^{2}(l-1).$
This proves the second statement and ends the proof of the lemma. ∎
Now we move to descent lemma in the local setup.
Lemma G.2.
Under Assumptions 2.1, 2.2(c), 2.3, 2.4, 2.5, at each iteration $t$ of Algorithm 7, the following estimation is valid:
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
Proof.
To begin with, we consider iterations, when Algorithm 7 performs aggregations, i.e. $t=t_{k\cdot l}$ for some $k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$. The update formula for such iterations is given by
$x^{t+1}=x_{i}^{t+1}=\sum\limits_{i=1}^{n}\left(\arg\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)\right]\right)\left(x_{i}^{t}-\gamma g_{i}^{t}\right),$
which leads to an upper bound on $\hat{f}(x^{t+1})$:
$\hat{f}(x^{t+1})~{}$ $~{}~{}\leqslant~{}\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x^{t}-\gamma g_{i}^{t}\right)\right]+\delta.$
Using this estimate and additional notation $\overline{x}^{t}=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}$ we proceed to the average per honest devices during this local round point estimate:
$\hat{f}\left(\overline{x}^{t+1}\right)$ $=$ $\hat{f}(x^{t+1})\leqslant~{}\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[\sum_{i=1}^{n}\omega_{i}\left(x_{i}^{t}-\gamma g_{i}^{t}\right)\right]+\delta$
$\leqslant$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta.$
Taking expectation, we obtain
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
$=$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\left\langle\nabla f\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
$-\left\langle\nabla f\left(\overline{x}^{t}\right)-\nabla\hat{f}\left(\overline{x}^{t}\right),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
[…]
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{\gamma}{2}\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\gamma\left\|\nabla\hat{f}\left(\overline{x}^{t}\right)-\nabla f_{1}\left(\overline{x}^{t}\right)\right\|^{2}+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
[…]
$+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}\left\|\nabla f\left(\overline{x}^{t}\right)-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\gamma\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\zeta(N)+\delta.$
Now we use Assumption 2.3 with the fact that $\mathbb{E}g_{i}^{t}=\nabla f_{i}(x^{t})$ and $\mathbb{E}\langle\nabla f_{i}(x^{t})-g_{i}^{t},\nabla f_{j}(x^{t})-g_{j}^{t}\rangle=0$ to bound $\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\left(g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right)\right\|^{2}\leqslant\frac{\sigma^{2}}{G}$. Taking expectation again, we move to
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma}{2}(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}$
$+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\gamma(1+3L\gamma)\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(1-3L\gamma\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}$
$+\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}+\gamma\mathbb{E}\left\|\nabla f_{1}\left(\overline{x}^{t}\right)-\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma\zeta(N)+\delta.$
Using Assumption 2.4 to bound $\left\|\nabla f_{i}(\overline{x}^{t})-\nabla f(\overline{x}^{t})\right\|^{2}\leqslant\delta_{1}+\delta_{2}\left\|\nabla f(\overline{x}^{t})\right\|^{2}$ for all $i\in\mathcal{G}$ we get
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$ (11)
$+\quad\frac{\gamma(1+3L\gamma)}{G}\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\nabla f_{i}(\overline{x}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}+L^{2}\gamma(1+3L\gamma)\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
Now we want to give the estimate for iterations $t\neq t_{k\cdot l}$. The update rule combined with (Lip) gives
$\hat{f}(\overline{x}^{t+1})$ $=$ $\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right]\leqslant\hat{f}\left[\frac{1}{G}\sum\limits_{i=1}^{n}x_{i}^{t}-\frac{\gamma}{G}\sum\limits_{i=1}^{n}g_{i}^{t}\right]+\delta.$
Mention, this estimate is coincide with the (G.2). Thus, proceeding analogically to the $t=t_{k\cdot l}$ case, we obtain (11). In that way, (11) delivers the result of the lemma. ∎
Now we pass to the main theorem of this section.
Theorem G.3.
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{8}$, 2.5, for solving the problem (1), for Algorithm 7 with $\gamma\leqslant\frac{1}{25(l-1)L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$$\leqslant$ $\frac{5\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+20L^{2}\gamma^{2}(l-1)\sigma^{2}+\frac{8L\gamma}{G}\sigma^{2}$
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N)+5\frac{\delta}{\gamma}.$
Proof.
To begin with, we combine the result of Lemma G.2 with result of Lemma G.1 to obtain
$\mathbb{E}\hat{f}\left(\overline{x}^{t+1}\right)$$\leqslant$ $\mathbb{E}\hat{f}\left(\overline{x}^{t}\right)-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{9\delta_{2}L\gamma^{2}(1+3L\gamma)}{2}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)}{2}$
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
Summing over all iterations and using $\sum\limits_{t=0}^{T-1}\sum\limits_{j=t_{k\cdot l}}^{t-1}\mathbb{E}\left\|\nabla f(\overline{x}^{j})\right\|^{2}\leqslant(l-1)\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}$,
$E\left[\hat{f}\left(x^{T}\right)-\hat{f}\left(x^{0}\right)\right]$$\leqslant$ $-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{9\delta_{2}L\gamma^{2}(1+3L\gamma)(l-1)}{2}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T+\delta T$
$=$ $-\frac{\gamma}{2}\left(1-4\delta_{2}-3L\gamma(1+2\delta_{2})-9L\gamma\delta_{2}(1+3L\gamma)(l-1)\right)\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T+\delta T.$
Now we want to estimate the coefficient before the $\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$ term. Let us take $\delta_{2}\leqslant\frac{1}{8}$ and $\gamma\leqslant\frac{1}{25(l-1)L}$ (it is the smallest $\gamma$ from all we choose before, thus, all previous transitions holds true). Thus,
$1-4\delta_{2}-3L\gamma(1+2\delta_{2})-9L\gamma\delta_{2}(1+3L\gamma)(l-1)$$\geqslant$ $\frac{1}{2}-\frac{15}{4}L\gamma-\frac{9}{8}L\gamma\left(1+\frac{3}{25(l-1)}\right)(l-1)$
$\overset{l\geqslant 2}{\geqslant}$ $\frac{1}{2}-\frac{15}{4}L\gamma-\frac{63}{50}L\gamma(l-1)$
$\geqslant$ $\frac{1}{2}-\frac{3}{20}-\frac{63}{1250}\geqslant\frac{1}{5}$
In that way,
$\frac{\gamma}{5}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$$\leqslant$ $\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]+4L^{2}\gamma^{3}(l-1)\sigma^{2}T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}$
$+\frac{5}{2}\gamma\delta_{1}T+6L\gamma^{2}\delta_{1}(l-1)T+\gamma\zeta(N)+\delta T,$
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$$\leqslant$ $\frac{5\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+20L^{2}\gamma^{2}(l-1)\sigma^{2}+\frac{8L\gamma}{G}\sigma^{2}$
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N)+5\frac{\delta}{\gamma},$
that ends the proof of the theorem. ∎
Remark G.4.
In this remark we want to explain why $\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}\right)\right\|^{2}$, that we choose as a criterion in Theorem G.3 is consistent. When we consider iterations, when Algorithm 7 performs aggregation, we assign weights of honest devices, i.e. devices, which send honest stochastic gradient $g$ with $\mathbb{E}[g]=\nabla f(x)$ equal to $\frac{1}{G}$, and weights of other devices (who acts as Byzantine) is equal to 00. And this layout gives an estimate that is an upper bound of true $\omega$ realization (G.2). In other words, we say, that if algorithm at each aggregation round take average of points from devices, which all previous local round act like honest it would not be better than for convergence, than real iteration of the algorithm. In such a way, we a not interesting in the points of devices, who perform even one Byzantine-like iteration in the local round. Now it is clear why we have to weaken the assumption about at least one honest device at each iteration, not necessary the same: we request at least one honest device at each local round, not necessary the same in different rounds.
Corollary G.5.
Under the assumptions of Theorem G.3, for solving the problem (1), after $T$ iterations of Algorithm 7 with $\gamma\leqslant\min\left\{\frac{1}{25(l-1)L},\frac{\sqrt{5\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]G}}{\sigma\sqrt{9LT}}\right\}$, the following holds:
[…]
We note that we do not provide the version for the first of our algorithms, BANT in this and the following sections. The reason is that its examination completely replicates the implementation of the technique discussed in this section for the proofs presented in Section D. We consider it unnecessary to repeat this procedure; however, we are confident in the practical applicability of the technique to both of our methods: BANT and AutoBANT.
Appendix H Partial participation
In the previous section, we discussed an important regime in distributed systems known as the local approach. Another widely used scenario in distributed learning is partial participation, which can be advantageous in various practical setups (Yang et al., 2021; Wang & Ji, 2022; Li et al., 2022). In this section, we extend our AutoBANT algorithm to support partial participation in federated learning. Below, we provide a formal description of the AutoBANT with Partial Participation method (Algorithm 8).
Algorithm 8 AutoBANT with Partial Participation
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$
2:  Parameters: Stepsize $\gamma>0$, error accuracy $\delta$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Define a set of active workers $\mathcal{W}(t)=\mathcal{G}(t)\cup\mathcal{B}(t)$$n(t)=|\mathcal{W}(t)|$
5:     Server sends $x^{t}$ to each worker from $\mathcal{W}(t)$
6:     for all workers $i\in\mathcal{W}(t)$ in parallel do
7:        Generate $\xi_{i}^{t}$ independently
8:        Compute stochastic gradient $g_{i}^{t}=g_{i}(x^{t},\xi_{i}^{t})$
9:        Send $g_{i}^{t}$ to server
10:     end for
11:     $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n(t)}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i\in\mathcal{W}(t)}\omega_{i}g_{i}^{t}\right)$
12:     $x^{t+1}=x^{t}-\gamma\sum\limits_{i\in\mathcal{W}(t)}\omega_{i}^{t}g_{i}^{t}$
13:  end for
14:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
The key change is that at each iteration we have a set $\mathcal{W}(t)$, which is a subset of the full list of accessible devices. In that way, we impose a stricter assumption compared to the regular training regime, namely, we require the presence of at least one honest device (including server) in each of the $\mathcal{W}(t)$ sets.
Now we present the main theorem of this section.
Theorem H.1.
Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}\leqslant 0.25$, 2.5, for solving the problem (1), after $T$ iteration of Algorithm 2 in partial participation scenario with $\gamma\leqslant\frac{1}{15L}$, it implies
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma},$
where $\widetilde{G}=\underset{t\leqslant T}{\min}G(t)$.
Proof.
The iterative update formula for $x^{t+1}$ is given by
$x^{t+1}$ $=$ $x^{t}-\gamma\sum_{i\in\mathcal{W}(t)}\left(\arg\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i\in\mathcal{W}(t)}\omega_{i}g_{i}^{t}\right]\right)g_{i}^{t},$
which leads to an upper bound on $\hat{f}(x^{t+1})$:
$\hat{f}(x^{t+1})$$\leqslant$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i\in\mathcal{W}(t)}\omega_{i}g_{i}^{t}\right]+\delta$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
Taking the expectation,
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
[…]
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x)\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f(x^{t})\right\|^{2}+\delta.$
Due to the fact that $\mathbb{E}g_{i}^{t}=\nabla f_{i}(x^{t})$ and $\mathbb{E}\langle\nabla f_{i}(x^{t})-g_{i}^{t},\nabla f_{j}(x^{t})-g_{j}^{t}\rangle=0$,
$\mathbb{E}\hat{f}(x^{t+1})$ (${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x)\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left(\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\frac{1}{(G(t))^{2}}\sum_{i\in\mathcal{G}(t)}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
[…]
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G(t)}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
$=$ $\mathbb{E}[\hat{f}(x^{t})]-\frac{\gamma}{2}\left[1-3L\gamma-(3+3L\gamma)\delta_{2}\right]\|\nabla f(x^{t})\|^{2}$
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G(t)}\sigma^{2}+\gamma\zeta(N)+\delta.$
We first fix $\delta_{2}\leqslant\frac{1}{12}$. Finally, by choosing $\gamma\leqslant\frac{1}{13L}\leqslant\frac{1}{12L(1+\delta_{2})}$, utilizing $G(t)\geqslant\underset{t\leqslant T}{\min}~{}G(t)=\widetilde{G}$ and summing over the iterations, we obtain the constraint in Theorem:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
Corollary H.2.
Under the assumptions of Theorem H.1, for solving the problem (1), after $T$ iterations with $\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]\widetilde{G}}}{\sigma\sqrt{3LT}}\right\}$, the following holds:
[…]
Remark H.3.
We adopted our method to the scenario of partial participation. It is natural that it requires stronger assumption: we need at least one honest client from that taking participation at each iteration, but not always the same.
Appendix I SimBANT
We also propose another method for dealing with Byzantine attacks - SimBANT. Here we use the concept of the trial function in a different way: the key is that we look not at the loss function $\hat{f}$, as in Algorithms 1, 2, but at the output of the model $m$ on the trial data $\hat{D}$. In particular, the calculation of trust scores $w_{i}$ is now based on how the outputs of the model parameters obtained on the server (a guaranteed honest device, without losing generality, we can assume that it has index 1) and on the device are similar to each other. The server is validated in pairs with each device and the trust scores are calculated for each device relative to the prediction classes based on how similar they are between the server and the device. And only on the basis of the trust scores, we give a weights to devices’ trained models and calculate the new state of the general model. Let us denote by $g^{t}$ the model that the server sends to the devices at step $t$, $g_{i}^{t},i\in\{2,\ldots,n\}$ devices’ trained models, $g_{1}^{t}$ the trained server model. Then the trust score for $i$-th device is function $\alpha_{i}\rightarrow sim(m(x^{t}-\gamma g_{i}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}))$, which measures the similarity of predictions of server and device models, where $x^{t}-\gamma g_{i}^{t}$ and $x^{t}-\gamma g_{1}^{t}$ are new parameters on the $i$-th device and the server, respectively, $sim$ is responsible for the measure of the similarity of the outputs. We do not provide the theory for this method but validate it in practice (see Section 5).
Algorithm 9 SimBANT
1:  Input: Starting point $x^{0}\in\mathbb{R}^{d}$
2:  Parameters: Stepsize $\gamma>0$, error accuracy $\delta$
3:  for $t=0,1,2,\ldots,T-1$ do
4:     Server sends $x^{t}$ to each worker
5:     for all workers $i=1,2,\ldots,n$ in parallel do
6:        Generate $\xi_{i}^{t}$ independently
7:        Compute stochastic gradient $g_{i}^{t}=g_{i}(x^{t},\xi_{i}^{t})$
8:        Send $g_{i}^{t}$ to server
9:     end for
10:     $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{sim(m(x^{t}-\gamma g_{i}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}{\sum\limits_{j=1}^{n}sim(m(x^{t}-\gamma g_{j}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}$
11:     $x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
12:  end for
13:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$